### v1

# Thermodynamics of quantum processes: An operational framework for free energy and reversible athermality

Abstract
We explore the thermodynamics of quantum processes (quantum channels) by axiomatically introducing the free energy for channels, defined via the quantum relative entropy with an absolutely thermal channel whose fixed output is in equilibrium with a thermal reservoir. This definition finds strong support through its operational interpretations in designated quantum information and thermodynamic tasks. We construct a resource theory of athermality for quantum processes, where free operations are Gibbs preserving superchannels and golden units are unitary channels with respect to absolutely thermal channel having fully degenerate output Hamiltonian. We exactly characterize the one-shot distillation and formation of quantum channels using hypothesis-testing and max-relative entropy with respect to the absolutely thermal channel. These rates converge asymptotically to the channel free energy (up to a multiplicative factor of half the inverse temperature), establishing its operational meaning and proving the asymptotic reversibility of the athermality. We show the direct relation between the resource theory of athermality and quantum information tasks such as private randomness and purity distillation and thermodynamic tasks of erasure and work extraction. Our work connects the core thermodynamic concepts of free energy, energy, entropy, and maximal extractable work of quantum processes to their information processing capabilities.
I Introduction
Quantum computation and information processing are fundamentally built upon quantum processes [1, 2, 3, 4, 5], which are the physical transformations and evolutions that quantum systems undergo. Studying the thermodynamic aspects of these systems and processes through an information-theoretic approach offers essential insights into the fundamental limitations on preparation, manipulation, and measurement of quantum resources, especially when faced with energetic constraints [6, 7, 8, 9, 10, 11]. Free energy stands as a core concept in thermodynamics, and its extension to quantum states provides insight to limitations on allowed transformations of states [12, 13, 14, 15, 16] and possible utilization for thermodynamic engines and circuits [17, 18, 19, 20].
Figure 1: A bipartite system with a non-interacting Hamiltonian $\widehat{H}_{R}+\widehat{H}_{A}$ is in state $\psi_{RA^{\prime}}$, in contact with a bath at inverse temperature $\beta$. A quantum channel $\mathcal{N}_{A^{\prime}\to A}$ acts locally on $A^{\prime}$ such that $\widehat{H}^{\rm int}_{RA}=0$. To extract work, we partially thermalize the channel output, $\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})\to\psi_{R}\otimes\gamma_{A}$. We consider partial thermalization composed of three steps: (1) Decoupling: $\mathcal{N}(\psi_{RA^{\prime}})=:\rho_{RA}\to\rho_{R}\otimes\rho_{A}$, (2) Quenching: The Hamiltonian is instantaneously changed from $\widehat{H}_{A}$ to $\beta^{-1}\ln\rho_{A}$ such that $\rho_{A}$ is an equilibrium state. (3) Isothermal quasistatic transformation of the Hamiltonian back to $\widehat{H}_{A}$. The extractable work from each step of the process is written in the blue color. The thermodynamic utility of the channel is quantified as the maximal extractable work via the partial thermalization process, see Theorem 6.
Concepts States Channels
Equilibration with bath Thermal state $\gamma^{\beta}$ Absolutely thermal channel $\mathcal{T}^{\beta}$
Energy $E(\rho)=\langle\widehat{H}\rangle_{\rho}$ $E[\mathcal{N}]=\sup_{\rho\in\operatorname{St}(A^{\prime})}\langle\widehat{H}\rangle_{\mathcal{N}(\rho)}$
(for channel output noninteracting with its reference)
Entropy $S(\rho)=-\operatorname{tr}[\rho\ln\rho]=-D(\rho\|\mathbbm{1})$ $S[\mathcal{N}]=-D[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]$
Thermodynamic free energy $F_{c}^{\beta}(\rho)=\beta^{-1}D(\rho\|\widehat{\gamma}^{\beta})$ $F_{c}^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$
Resource-theoretic free energy $F^{\beta}(\rho)=F^{\beta}_{c}(\rho)-F^{\beta}_{c}(\gamma^{\beta})=\beta^{-1}D(\rho\|{\gamma}^{\beta})$ $F^{\beta}[\mathcal{N}]=F_{c}^{\beta}[\mathcal{N}]-F_{c}^{\beta}[\mathcal{T}^{\beta}]=\beta^{-1}D[\mathcal{N}\|{\mathcal{T}}^{\beta}]$
Minimal free energy Thermal state $\gamma^{\beta}$ Absolutely thermal channel $\mathcal{T}^{\beta}$
Maximal extractable work $W^{\rm ext}_{\rho\to\gamma}=F^{\beta}_{c}(\rho)-F_{c}^{\beta}(\gamma^{\beta})=F^{\beta}(\rho)$ $W^{\rm ext}_{\mathcal{N}\to\mathcal{T}^{\beta}}=F_{c}^{\beta}[\mathcal{N}]-F_{c}^{\beta}[\mathcal{T}^{\beta}]=F^{\beta}[\mathcal{N}]$
Helmholtz equation $F^{\beta}_{c}(\rho)=E(\rho)-\beta^{-1}S(\rho)$ $F^{\beta}_{c}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}]$
(for channel output noninteracting with its reference; inequality saturates for replacer channels)
Table 1: We summarize and compare mathematical expressions for elementary thermodynamic concepts for quantum states vs channels. We consider thermal reservoir to be at inverse temperature $\beta$ and $\widehat{H}$ denotes the Hamiltonian of the designated system. $\widehat{\gamma}^{\beta}$ is a thermal (Gibbs) operator and $\gamma^{\beta}$ is a thermal state. $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\omega_{A}$, $\mathcal{T}^{\beta}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\gamma^{\beta}_{A}$, $\widehat{\mathcal{T}}^{\beta}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\widehat{\gamma}^{\beta}$. $D(\rho\|\sigma):=\lim_{\varepsilon\to 0^{+}}\operatorname{tr}[\rho(\ln\rho-\ln(\sigma+\varepsilon\mathbbm{1}))]$ is (Umegaki) relative entropy between states $\rho,\sigma$ and $D[\mathcal{N}\|\mathcal{M}]:=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D(\operatorname{id}_{R}\otimes\mathcal{N}(\psi_{RA^{\prime}})\|\operatorname{id}_{R}\otimes\mathcal{M}(\psi_{RA^{\prime}}))$ is relative entropy between channels $\mathcal{N}_{A^{\prime}\to A},\mathcal{M}_{A^{\prime}\to A}$. We assume $\rho$ is a density operator of system $A^{\prime}$ and $\mathcal{N}$ is a quantum channel from $A^{\prime}\to A$. The content for quantum states are known in literature [12, 21, 22, 23], prior to our current work. The expression of the entropy of quantum channels is from [24, 25]. See Section I.2 for the outline of the paper.
Motivated by fundamental interests and technological imperative of miniaturization, we explore the thermodynamics of quantum processes and their implications for information processing capabilities. Quantum processes are formally described by quantum channels, which are completely positive and trace-preserving (CPTP) maps. Quantum channels capture all physically realizable transformations of quantum states, including quantum measurements and even states themselves, when regarded as preparation channels. Although the action of a channel can be inferred from how it transforms input states, the information-theoretic properties established for quantum states do not always extend directly to channels. Operational tasks such as channel discrimination and resource conversion reveal richer and more intricate structures at the level of quantum channels [26, 27, 28, 29, 30]. However, many information-theoretic results for channels naturally recover the corresponding results for states as special cases, since quantum states can be regarded as preparation channels.
Our objective is the quantification and manipulation of the thermodynamic resources intrinsic to quantum channels [19, 31]. The thermodynamic characterization of general quantum channels necessitates the development of new conceptual frameworks. In this work, we use an axiomatic approach [21, 12, 32] to define a notion of free energy for quantum channels, which enables us to formulate a resource theory of thermodynamic processes where this notion of free energy acquires a clear operational meaning. This framework also enables the analysis of how core thermodynamic quantities such as free energy, entropy, and energy appear at the level of quantum processes, and how they constrain or enable information processing tasks. This unified view establishes formal links between the thermodynamics of quantum channels and operational tasks such as randomness extraction, purity distillation, and work extraction [19, 33, 34, 35, 36, 37, 38].
The process of thermalization, which describes non-unitary evolution toward equilibrium with an environment (thermal bath) [39, 40, 41, 21, 42, 43], serves as a concrete example of a thermodynamic quantum process. It is characterized by the irreversible loss of coherence and the emergence of thermal states that minimize the system’s free energy [44, 45, 46, 31]. In this context, the concepts of free energy and the resource theory of athermality for quantum channels provide a rigorous framework for quantifying thermodynamic costs and understanding the fundamental limitations of quantum operations, thereby offering a clear connection between quantum information theory and foundational thermodynamic principles.
I.1 Main Results
The concept of the free energy is fundamental and instrumental in thermodynamics. We extend this notion to quantum processes which describe physical transformations of quantum systems in quantum theory. We leverage axiomatic and information-theoretic approaches to define meaningful free energy functions of quantum channels (Theorem 1) that find operational interpretations in quantum information processing and thermodynamical tasks. We establish relations between the elementary thermodynamic quantities like free energy, energy, entropy, and operational quantity like extractable work to shed light on the thermodynamics of quantum processes (see Table 1).
We define the resource-theoretic generalized free energy functions $\mathbf{F}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ as the appropriate generalized channel divergence functions, $\mathbf{F}^{\beta}[\mathcal{N}]:=\beta^{-1}\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$ [25, 30], between the channel $\mathcal{N}$ and an absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$. The channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ thermalizes all its inputs such that the output state is a thermal state $\gamma^{\beta}_{A}$ in equilibrium with the bath at an inverse temperature $\beta>0$, $\mathcal{T}^{\beta}(\cdot)=\operatorname{tr}[\cdot]\gamma^{\beta}$. Some appropriate generalized channel divergences for our purpose are shown to be the Umegaki relative entropy and max-relative entropy. We also discuss smoothed versions of free energy functions derived from these channel divergences and the $\varepsilon$-hypothesis-testing relative entropy.
We develop the dynamical resource theory of athermality for (square) quantum channels. The absolutely thermal channel $\mathcal{T}^{\beta}$ is considered a free object, where $\mathcal{T}^{\beta}$ depends on the inverse temperature of the bath and the Hamiltonian of the output system under consideration. This is a natural assumption because of the inherent thermalization process that equilibrates quantum systems with the bath. Even near isolation of quantum systems from the bath for short time requires engineering and interventions, and perfect closed quantum dynamics are idealistic. The free operations are Gibbs preserving superchannels $\Theta^{\beta}$ which map absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ to absolutely thermal channel $\mathcal{T}^{\beta}_{B^{\prime}\to B}$. A quantum channel $\mathcal{N}$ is athermal if it isn’t asbolutely thermal $\mathcal{T}^{\beta}$ and is deemed resourceful. For clarity, we use $(\mathcal{N},\mathcal{T}^{\beta})$ to discuss the thermodynamic resourcefulness of $\mathcal{N}$ with respect to $\mathcal{T}^{\beta}$ [31].
The resource-theoretic (and thermodynamic) free energy and max-free energy are maximum for a square quantum channel if and only if it’s a unitary channel (Theorem 2). Among all possible Hamiltonians, the free energy of a unitary channel $\mathcal{U}_{A^{\prime}\to A}$ is least if the output Hamiltonian is fully degenerate, $\widehat{H}_{A}=c\mathbbm{1}_{A}$ for $c\in\mathbbm{R}$. With certain physical motivation, we advocate that a golden unit of the dynamical resource theory of athermality is the resource channel $(\operatorname{id}_{m},\mathcal{R}^{\gamma})$, where ${\mathcal{N}}_{m}$ denotes that input and output Hilbert spaces of channel $\mathcal{N}$ are $m$-dimensional and $\mathcal{R}^{\pi}$ is the uniformly mixing channel that only outputs the maximally mixed state $\pi$. All unitary resource channels $(\mathcal{U}_{m},\mathcal{R}^{\pi})$ are equivalent (equally resourceful) to $(\operatorname{id}_{m},\mathcal{R}^{\pi})$. In fact, the interconversion between $(\mathcal{U},\mathcal{T}^{\beta})\leftrightarrow(\mathcal{V},\mathcal{T}^{\beta})$ is possible via Gibbs preserving superchannels (free operations) for any pair of unitary channels $\mathcal{U}$ and $\mathcal{V}$.
We find that the single-shot athermality distillation $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and formation $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, up to allowed error $\varepsilon\in(0,1)$ under the action of Gibbs preserving superchannels, are $\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$ and $\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$, respectively (Theorem 3). We have $F^{\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$, which provide exact operational meaning to the resource-theoretic hypothesis-testing free energy and the max-free energy of a quantum channel. The asymptotic athermality distillation and formation rates of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, irrespective of adaptive or nonadaptive strategies, are equal to $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ (Theorem 4). The resource-theoretic free energy of a channel $\mathcal{N}$, $F^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, and thus its asymptotic athermality distillation and formation rates are both equal to $\frac{\beta}{2}F^{\beta}[\mathcal{N}]$, providing the exact operational interpretation to the resource-theoretic free energy of a quantum channel. Both asymptotic athermality distillation and formation rates of a quantum channel are the same, implying that the athermality of quantum channels under Gibbs preserving superchannels is (asymptotically) reversible.
We get the resource theory of purity for quantum channels as a special case in the limit $\beta\to 0^{+}$ or Hamiltonian of the channel output is trivial. We discuss the dependence of the free energy of a quantum channel on quantum transmission ability of the channel and free energy of its output (Proposition 5). We elucidate on the novel relations between the free energy of a quantum channel and its energy and entropy. The thermodynamic free energy $F_{c}[\mathcal{N}]:=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$, where $\widehat{\mathcal{T}}^{\beta}$ is a replacer (or preparation) map that always outputs thermal operator $\widehat{\gamma}^{\beta}$ (an unnormalized thermal state), of a quantum channel $\mathcal{N}$ and its entropy $S[\mathcal{N}]=-D[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]$ [24, 25], where $\mathcal{R}^{\mathbbm{1}}(\cdot)=\operatorname{tr}[\cdot]\mathbbm{1}$, and thermodynamic energy $E[\mathcal{N}]$ satisfy the relation $F_{c}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}]$ when the channel output is noninteracting with its reference (Theorem 5).
We prove that the resource-theoretic free energy of a quantum channel $\mathcal{N}$ is the difference of the thermodynamic free energy of the channel and the absolutely thermal channel, $F[\mathcal{N}]=F_{c}[\mathcal{N}]-F_{c}[\mathcal{T}^{\beta}]$, and has operational interpretation as the optimal (maximum) thermodynamic utility of the channel for maximal extractable work (Theorem 6, see Fig. 1). We believe our framework, results, and discussions will provide deeper insights on thermodynamical aspects of quantum processes and interconnectedness with their quantum information processing capabilities.
I.2 Outline
The paper is organized as follows. We introduce standard notations, definitions, and facts in Section II that are used to derive main results and discussion surrounding them. In Section III, we employ axiomatic approach to define the resource-theoretic free energy functions of a quantum channel and provide reasoning behind axioms. We discuss some notable definitions, namely, the (resource-theoretic) free energy, max-free energy, and hypothesis-testing free energy of quantum channels and prove some of their important properties. We introduce the resource theory of athermality for square quantum channels in Section IV and prove that the resource theory of athermality under Gibbs preserving superchannels is asymptotically reversible. We inspect quantum transmission capabilities of quantum channels and its free energy in Section V. We discuss relation between the resource theory of athermality with private randomness distillation and dynamical resource theory of purity. In Section VI, we delve deeper into the relations between the resource-theoretic free energy, thermodynamic free energy, entropy, and energy of quantum channels. We discuss operational meaning of the resource-theoretic free energy of a quantum channel with its thermodynamic utility for the maximal extractable work. We finally conclude in Section VII. The detailed proofs of several results are discussed in Appendix.
II Preliminaries
We consider quantum systems associated with separable Hilbert spaces with tensor-product structure. For a bipartite state $\rho_{AB}$, $\rho_{A}=\operatorname{tr}_{B}(\rho_{AB})$. The mean energy of a quantum state $\rho_{A}$ associated with Hamiltonian $\widehat{H}_{A}$ is $\langle\widehat{H}\rangle_{\rho}:=\operatorname{tr}[\widehat{H}_{A}\rho_{A}]$. Let $\operatorname{St}(A)$ and $\operatorname{Pos}(A)$ denote the set of all density operators (quantum states) and the set of all positive semidefinite operators on $A$, respectively. Let $\operatorname{Ch}(A^{\prime},A)$ denote the set of all quantum channels from $A^{\prime}$ to $A$. Let $\operatorname{SCh}((A^{\prime},A),(B^{\prime},B))$ denote the set of all quantum superchannels mapping $\operatorname{Ch}(A^{\prime},A)$ to $\operatorname{Ch}(B^{\prime},B)$. Any quantum superchannel can be written as concatenation of preprocessing and postprocessing quantum channels [47], i.e., for any $\Theta\in\operatorname{SCh}((A^{\prime},A),(B^{\prime},B))$, there exists preprocessing channel $\mathcal{P}\in\operatorname{Ch}(B^{\prime},RA^{\prime})$ and postprocessing channel $\mathcal{Q}\in\operatorname{Ch}(RA,B)$ such that $\Theta(\mathcal{N})=\mathcal{Q}\circ\mathcal{N}\circ\mathcal{P}$. A unitary superchannel is composed of unitary postprocessing and preprocessing channels.
[…]
Generalized state and channel divergences.— Let $\mathbf{D}(\cdot\|\cdot)$ be the generalized state divergence, i.e., for an arbitrary pair of quantum states $\rho,\sigma$ and an arbitrary quantum channel $\mathcal{N}$, we have $\mathbf{D}(\rho\|\sigma)\geq\mathbf{D}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma))$. The generalized channel divergence $\bf{D}[\cdot\|\cdot]$ is derived from the generalized state divergence: For any two quantum channels $\mathcal{N},\mathcal{M}\in\operatorname{Ch}(A^{\prime}A)$ and the identity channel $\operatorname{id}$,
$\mathbf{D}[\mathcal{N}\|\mathcal{M}]:=\sup_{\psi\in\mathrm{St}(RA^{\prime})}\mathbf{D}(\operatorname{id}_{R}\otimes\mathcal{N}(\psi)\|\operatorname{id}_{R}\otimes\mathcal{M}(\psi)),$ (1)
where it suffices to optimize over pure states $\psi_{RA^{\prime}}$ [48]. Some examples of the generalized state divergences [49, 50, 51, 52]: For a state $\rho_{A}$ and a positive semidefinite operator $\sigma_{A}$, whenever $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$ (otherwise $\infty$) (a) the quantum relative entropy $D(\rho\|\sigma):=\operatorname{tr}\left[\rho(\ln\rho-\ln\sigma)\right]$, (b) the max-relative entropy $D_{\infty}(\rho\|\sigma):=\ln\norm{\sigma^{-1/2}\rho\,\sigma^{-1/2}}_{\infty}$, (c) the sandwiched Rényi relative entropy ${D}_{\alpha}(\rho\|\sigma):=\frac{1}{\alpha-1}\ln\left(\operatorname{Tr}\left[\left(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\,\sigma^{\frac{1-\alpha}{2\alpha}}\right)^{\alpha}\right]\right)$ for $\alpha\in[\frac{1}{2},1)\cup(1,\infty)$, (d) the $\varepsilon$-hypothesis-testing relative entropy $D^{\varepsilon}_{\mathrm{H}}(\rho\|\sigma):=-\ln\inf_{0\leq\Lambda\leq\mathbbm{1}}\left\{\operatorname{tr}[\Lambda\sigma]\,:\,\operatorname{tr}[\Lambda\rho]\geq 1-\varepsilon\right\}$, (e) the purified distance $\mathrm{P}(\rho,\sigma):=\sqrt{1-F(\rho,\sigma)}$. $\ln$ denotes natural logarithm. We have $\lim_{\alpha\to 1}D_{\alpha}(\rho\|\sigma)=D(\rho\|\sigma)$ and $\lim_{\alpha\to\infty}D_{\alpha}(\rho\|\sigma)=D_{\infty}(\rho\|\sigma)$.
The generalized channel divergences are monotonically nonincreasing under the action of quantum superchannels [30, Theorem 1]. If the generalized state divergence is nonnegative, then the corresponding channel divergence is also nonnegative. If the generalized state divergence is faithful, i.e., minimum if and only the states are equal, then the corresponding channel divergence is also faithful, i.e., attains minimum if and only if the channels are equal.
An $\varepsilon$-ball around a state $\rho_{A}$, for $\varepsilon\in[0,1]$, is defined as $\mathcal{B}^{\varepsilon}(\rho_{A})=\{\sigma_{A}:\penalty 10000\ \sigma_{A}\geq 0,\operatorname{tr}(\sigma_{A})\leq 1,P(\rho_{A},\sigma_{A})\leq\varepsilon\}$. An $\varepsilon$-ball around a channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as $\mathcal{B}^{\varepsilon}[\mathcal{N}]=\{\mathcal{M}\in\operatorname{Ch}(A^{\prime},A):P[\mathcal{N},\mathcal{M}]\leq\varepsilon\}$. For $\rho\in\operatorname{St}(A)$ and $\sigma\in\operatorname{Pos}(A)$, $D_{\infty}^{\epsilon}(\rho\|\sigma)=\inf_{\omega\in\mathcal{B}^{\varepsilon}(\rho)}D_{\infty}(\omega\|\sigma)$. $D^{\varepsilon}_{\infty}(\cdot\|\cdot)$ is a generalized state divergence and $D^{\varepsilon}_{\infty}[\cdot\|\cdot]$ is a generalized channel divergence.
The generalized entropy function $\mathbf{S}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as $\mathbf{S}[\mathcal{N}]:=-\mathbf{D}[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]$ [24, 25], where a completely positive map $\mathcal{R}^{\omega}_{A^{\prime}\to A}$ denotes the replacer (or preparation) map that outputs $\omega\in\operatorname{Pos}(A)$, $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\operatorname{tr}[\rho_{A^{\prime}}]\omega_{A}$ for all $\rho_{A^{\prime}}\in\operatorname{Pos}(A^{\prime})$. For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ with $|A|<\infty$, $S_{\alpha}[\mathcal{N}]:=\ln|A|-D_{\alpha}[\mathcal{N}\|\mathcal{R}^{\pi}]$ [24], where $\alpha\in[\frac{1}{2},\infty]$ and $\alpha=1$ is $\lim_{\alpha\to 1}$. The (von Neumann) entropy of a quantum state $\rho_{A}$ is $S(\rho_{A}):=S(A)_{\rho}:=-D(\rho\|\mathbbm{1})=-\operatorname{tr}[\rho\ln\rho]$. $\mathcal{R}^{\pi}$ is called uniformly mixing or completely depolarizing channel as it always outputs the maximally mixed state $\pi(\propto\mathbbm{1})$.
III Dynamical free energy
The dynamical free energy, i.e., the free energy of quantum channels, should satisfy certain desirable properties analogous to the free energy of quantum states. It becomes operationally meaningful if it is related to the athermality distillation or formation tasks under relevant free operations. It is also preferable that the channel fee energy finds operational meaning in a thermodynamic task like work extraction.
Gibbs preservation.— For an inverse temperature $\beta\in\mathbbm{R}^{+}$, the absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ is a replacer (or preparation) channel that only outputs the thermal state $\gamma^{\beta}_{A}=\exp(-\beta\widehat{H}_{A})/{Z^{\beta}}$, where $\widehat{H}_{A}$ denotes the Hamiltonian of $A$ and $Z_{A}^{\beta}:=\operatorname{tr}[\exp(-\beta\widehat{H}_{A})]$ denotes the associated partition function [53]. The thermal state $\gamma^{\beta}_{A}$ remains invariant under the translation of the Hamiltonian, $\widehat{H}_{A}\to\widehat{H}_{A}+c\mathbbm{1}_{A}$ for $c\in\mathbbm{R}$. This implies that the absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ also remains invariant under the translation of the output Hamiltonian, $\widehat{H}_{A}\to\widehat{H}_{A}+c\mathbbm{1}_{A}$. We refer to a superchannel $\Theta^{\beta}\in\operatorname{SCh}((A^{\prime},A),(B^{\prime},B))$ as Gibbs-subpreserving if $\Theta^{\beta}(\mathcal{T}^{\beta}_{A^{\prime}\to A})\leq\mathcal{T}^{\beta}_{B^{\prime}\to B}$, i.e.,
$\forall\penalty 10000\ \psi\in\operatorname{St}(RB^{\prime}),\quad\Theta^{\beta}(\mathcal{T}^{\beta}_{A^{\prime}\to A})(\psi_{RB^{\prime}})\leq\psi_{R}\otimes\gamma^{\beta}_{B},$ (2)
and Gibbs preserving if $\Theta^{\beta}(\mathcal{T}^{\beta}_{A^{\prime}\to A})=\mathcal{T}^{\beta}_{B^{\prime}\to B}$.
For $|A|<\infty$ and $\widehat{H}_{A}=c\mathbbm{1}_{A}$, the thermal state is the maximally mixed state, $\gamma^{\beta}=\pi_{A}$ for all $\beta\in\mathbbm{R}^{+}$; $\gamma^{\beta}_{A}=\pi_{A}$ is true also for arbitrary $\widehat{H}_{A}$ when $\beta\to 0^{+}$. The entropy $S[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ with the fixed mean energy, $\sup_{\rho\in\operatorname{St}(A^{\prime})}\langle\widehat{H}\rangle_{\mathcal{N}(\rho)}=E$, is maximum if and only if the channel is absolutely thermalizing $\mathcal{T}^{\beta}$ with the thermal state of mean energy $\langle\widehat{H}\rangle_{\gamma^{\beta}}=E$ [54, 55, 31].
Taking a cue from the discussions in [31, 25, 30], we introduce the notion of the free energy of a quantum channel, which is pivotal to build the framework for a thermodynamic resource theory of quantum channels.
Thermodynamic and resource-theoretic axioms.— Let $f[\mathcal{N}]$ denote a physically motivated free energy function of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ in a resource-theoretic paradigm. The function $f:\operatorname{Ch}(A^{\prime},A)\to\mathbbm{R}^{+}$ should satisfy the following properties, (A1-A3) essentially and also (A4-A6) preferably, for a fixed inverse temperature $\beta$:
  1. (A1)
    Monotonically nonincreasing under the action of an Gibbs preserving superchannel $\Theta^{\beta}$, $\Theta^{\beta}(\mathcal{T}^{\beta}_{A^{\prime}\to A})=\mathcal{T}^{\beta}_{B^{\prime}\to B}$, $f[\Theta^{\beta}(\mathcal{N})]\leq f[\mathcal{N}]$.
  2. (A2)
    Reduction to the free energy function of the output state for a replacer channel $\mathcal{R}^{\omega}$, $\mathcal{R}^{\omega}(\rho_{A^{\prime}}):=\omega\in\operatorname{St}(A)$ for all input states $\rho_{A^{\prime}}$, $f[\mathcal{R}^{\omega}]=f(\omega)$.
  3. (A3)
    Minimum is attained if and only if the channel is absolutely thermal.
  4. (A4)
[…]
  5. (A5)
    Additive under tensor-product channels $\mathcal{N}_{A_{1}^{\prime}\to A_{1}}\otimes\mathcal{M}_{A_{2}^{\prime}\to A_{2}}$ with $\widehat{H}_{A_{1}A_{2}}=\widehat{H}_{A_{1}}\otimes\mathbbm{1}_{A_{2}}+\mathbbm{1}_{A_{1}}\otimes\widehat{H}_{A_{2}}$, $f[\mathcal{N}\otimes\mathcal{M}]=f[\mathcal{N}]+f[\mathcal{M}]$.
  6. (A6)
    Convexity, $f[p\mathcal{N}+(1-p)\mathcal{M}]\leq pf[\mathcal{N}]+(1-p)f[\mathcal{M}]$ for $p\in[0,1]$, where $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{A^{\prime}\to A}$ are quantum channels.
[…]
To construct a resource-theoretic framework and coin the dynamical free energy, it is important that the free channel is the absolutely thermal channel with its output in thermal equilirbium with bath. That is, we need to fix the inverse temperature corresponding to the bath. The free energy function of a quantum channel should capture how distinguishable it is from the absolutely thermal channel [31], and should therefore be monotonically nonincreasing under the actions of Gibbs preserving superchannels. Moreover, the thermodynamic resource theory of quantum states should be recoverable from the thermodynamic resource theory of quantum channels, since quantum states can be seen as preparation channels. For the free energy function of quantum channels to be physically meaningful, it should satisfy essential properties such as continuity and additivity, and preferably convexity (or quasi-convexity), which are well-known for free energy functions of states. The axiomatic properties (A1–A6) are designed to capture these requirements.
Definition 1 (Generalized free energy).
For a fixed $\beta$, the (resource-theoretic) generalized free energy function $\mathbf{F}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as the generalized divergence between the channel $\mathcal{N}$ and the absolutely thermal channel $\mathcal{T}^{\beta}$,
$\mathbf{F}^{\beta}[\mathcal{N}]:={\beta}^{-1}\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (3)
$\mathcal{T}^{\beta}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\gamma^{\beta}_{A}$ for all input states $\rho_{A^{\prime}}$.
The generalized free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ can also be expressed as
$\mathbf{F}^{\beta}[\mathcal{N}]=\frac{1}{\beta}\sup_{\psi\in\operatorname{St}(RA^{\prime})}\mathbf{D}(\operatorname{id}_{R}\otimes\mathcal{N}(\psi_{RA^{\prime}})\|\psi_{R}\otimes\gamma^{\beta}_{A}),$ (4)
where it suffices to optimize over pure states $\psi_{RA^{\prime}}$. The absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ reflects the thermal state of the system $A$ that is in equilibrium with the bath; it entails both, the inverse temperature $\beta$ associated with the bath and the Hamiltonian $\widehat{H}_{A}$ of the output system.
In the definition of the generalized free energy $\mathbf{F}^{\beta}[\mathcal{N}]$, Eqs.(3) and (4), we could possibly fix the Hamiltonian of the reference $R$ to the channel $\mathcal{N}$ as per need. For instance, we can always consider $\widehat{H}_{R}=0$ with the total Hamiltonian of the reference and output to be $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes\widehat{H}_{A}$, if we want the mean energy of the process $\operatorname{id}_{R}\otimes\mathcal{N}$ and $\mathcal{N}$ to be the same:
$\sup_{\psi\in\operatorname{St}(RA^{\prime})}\operatorname{tr}[\mathbbm{1}_{R}\otimes\widehat{H}_{A}(\operatorname{id}_{R}\otimes\mathcal{N}(\psi_{RA^{\prime}}))]$
$\qquad\qquad=\sup_{\phi\in\operatorname{St}(A^{\prime})}\operatorname{tr}[\widehat{H}_{A}\mathcal{N}(\phi_{A^{\prime}})].$ (5)
See Section VI for a detailed description on the relation between the free energy, entropy, and energy of a quantum channel.
For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, (a) the free energy $F^{\beta}[\mathcal{N}]:=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, (b) the max-free energy $F^{\beta}_{\infty}[\mathcal{N}]:=\beta^{-1}D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]$, (c) the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]$, (d) the $\varepsilon$-hypothesis-testing free energy ${F}^{\beta,\varepsilon}_{\mathrm{H}}[\mathcal{N}]:=\beta^{-1}D^{\varepsilon}_{\mathrm{H}}[\mathcal{N}\|\mathcal{T}^{\beta}]$ for $\varepsilon\in[0,1]$, (e) ${F}^{\beta,\varepsilon}_{\infty}[\mathcal{N}]:=\beta^{-1}\inf_{\mathcal{M}\in\mathcal{B}^{\varepsilon}[\mathcal{N}]}D^{\varepsilon}_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]$ for $\varepsilon\in[0,1]$.
For tensor-product quantum channels $\mathcal{N}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}$, we say that the channels $\mathcal{N}$ and $\mathcal{M}$ are noninteracting when the total Hamiltonian of the outputs is of the form $\widehat{H}_{AB}=\widehat{H}_{A}\otimes\mathbbm{1}_{B}+\mathbbm{1}_{A}\otimes\widehat{H}_{B}$, i.e., the interaction Hamiltonian $\widehat{H}^{\rm int}_{AB}=0$. When the total Hamiltonian of the output is $\widehat{H}_{AB}=\widehat{H}_{A}\otimes\mathbbm{1}_{B}+\mathbbm{1}_{A}\otimes\widehat{H}_{B}$, the absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}B^{\prime}\to AB}$ can be written as the tensor-product of local absolutely thermal channels, $\mathcal{T}^{\beta}_{A^{\prime}B^{\prime}\to AB}=\mathcal{T}^{\beta}_{A^{\prime}\to A}\otimes\mathcal{T}^{\beta}_{B^{\prime}\to B}$.
We observe that the free energy of a quantum channel appears as the maximal conditional free energy associated with the channel, analogous to the channel entropy [56, 24].
Proposition 1.
The free energy $F^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is equal to the maximum channel output free energy conditioned on its reference,
$F^{\beta}[\mathcal{N}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[F^{\beta}(\operatorname{id}_{R}\otimes\mathcal{N}(\psi_{RA^{\prime}}))-F^{\beta}(\psi_{R})\right],$ (6)
where it suffices to optimize over pure states $\psi_{RA^{\prime}}$, $|R|=|A^{\prime}|$.
Proof.
The proof follows from the definition of the free energy of a quantum channel along with the facts that $F(\rho)=\beta^{-1}D(\rho\|\gamma^{\beta})$ and
$\sup_{\psi\in\operatorname{St}(RA^{\prime})}D(\mathcal{N}(\psi_{RA^{\prime}})\|\psi_{R}\otimes\gamma^{\beta}_{A})$
$=$ $\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[D(\mathcal{N}(\psi_{RA^{\prime}})\|\gamma^{\beta}_{R}\otimes\gamma^{\beta}_{A})-D(\psi_{R}\|\gamma^{\beta}_{R})\right],$ (7)
where we used the relation:
$\operatorname{tr}[\rho_{RA}\ln(\sigma_{R}\otimes\omega_{A})]=\operatorname{tr}[\rho_{R}\ln\sigma_{R}]+\operatorname{tr}[\rho_{A}\ln\omega_{A}].$ (8)
Proposition 2.
The max-free energy $F^{\beta}_{\infty}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is equal to
$F^{\beta}_{\infty}[\mathcal{N}]=\beta^{-1}D_{\infty}(\Phi^{\mathcal{N}}_{RA}\|\pi_{A}\otimes\gamma^{\beta}_{A}).$ (9)
For total Hamiltonian of the reference to the channel and its output $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$, where $c\in\mathbbm{R}$, the max-free energy of the channel is equal to the max-free energy of its Choi state,
$F^{\beta}_{\infty}[\mathcal{N}]=F^{\beta}_{\infty}(\Phi^{\mathcal{N}}_{RA}).$ (10)
Proof.
The proof of Eq. (9) follows from the fact that $D_{\infty}[\mathcal{N}\|\mathcal{M}]=D_{\infty}(\Phi^{\mathcal{N}}_{RA}\|\Phi^{\mathcal{M}}_{RA})$ [48],
$D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]$ $=D_{\infty}({\mathcal{N}}(\Phi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\Phi_{RA^{\prime}}))$ (11)
$=D_{\infty}(\Phi^{\mathcal{N}}_{RA}\|\pi_{A}\otimes\gamma^{\beta}_{A}).$ (12)
The proof of Eq. (10) as follows from the above. We notice that for $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$ (or, $\widehat{H}_{R}=c\mathbbm{1}_{R}$), $\pi_{R}\otimes\gamma^{\beta}_{A}=\gamma^{\beta}_{R}\otimes\gamma^{\beta}_{A}=\gamma^{\beta}_{RA}$. Then,
$F^{\beta}_{\infty}[\mathcal{N}]$ $=\beta^{-1}D_{\infty}(\Phi^{\mathcal{N}}_{RA}\|\gamma^{\beta}_{RA})=F^{\beta}_{\infty}(\Phi^{\mathcal{N}}_{RA}).$ (13)
III.1 Properties of the dynamical free energy
All the generalized free energy functions $\mathbf{F}^{\beta}[\mathcal{N}]$ satisfy (A1) and (A2) readily, whenever $\beta^{-1}\mathbf{D}(\rho\|\gamma^{\beta})$ represents some resource-theoretic free energy function $f(\rho)$ for states $\rho$ [12, 23]. The generalized free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ remains invariant under the action of Gibbs preserving unitary superchannels. Faithfulness (A3) is satisfied by the generalized free energy whose generalized channel divergence is faithful. For all $\alpha\in[\frac{1}{2},\infty)$, using $\lim_{\alpha\to 1}$ as $\alpha=1$, the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]\geq 0$ for an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$, and $F^{\beta}_{\alpha}[\mathcal{N}]=0$ iff $\mathcal{N}=\mathcal{T}^{\beta}$. For $\alpha\geq 1$ and $n\in\mathbbm{N}$, $F_{\alpha}^{\beta}[\mathcal{N}^{\otimes n}]=nF^{\beta}_{\alpha}[\mathcal{N}]$ if the outputs of the parallel uses of the channel $\mathcal{N}$ are noninteracting (i.e., the interaction Hamiltonian between the channel outputs is zero). These assertions follow from [30, Theorem 1] and [57, Theorem 4.2].
Theorem 1.
Let $\mathcal{N}_{A^{\prime}\to A}$ be an arbitrary quantum channel. The free energy $F^{\beta}[\mathcal{N}]$ satisfies all the thermodynamical (resource-theoretic) axiomatic properties (A1-A6).
For $\alpha\in(1,\infty)$, the sandwiched Rényi free energy $F_{\alpha}^{\beta}[\mathcal{N}]$ satisfies (A1-A5) and quasi-convexity, i.e., for quantum channel $\mathcal{N}=\sum_{x}p_{x}\mathcal{N}^{x}$ where $\{p_{x}\}_{x}$ is some probability distribution, $F_{\alpha}^{\beta}[\mathcal{N}]\leq\max_{x}F_{\alpha}^{\beta}[\mathcal{N}^{x}]$.
Proof.
We provide a detailed proof in Appendix A.1. ∎
Theorem 2.
For $\alpha\in\{1,\infty\}$ and $|A^{\prime}|=|A|$, the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is maximum if and only if $\mathcal{N}$ is a unitary channel. For all unitary channels $\mathcal{U}_{A^{\prime}\to A}$, $F^{\beta}_{\alpha}[\mathcal{U}]=F^{\beta}_{\alpha}[\operatorname{id}_{A^{\prime}\to A}]$.
Details of the proof are in Appendix A.3.
Proposition 3.
The max-free energy $F^{\beta}_{\infty}[\mathcal{U}]$ of a unitary quantum channel $\mathcal{U}_{A^{\prime}\to A}$ is
$F^{\beta}_{\infty}[\mathcal{U}]=\beta^{-1}\ln\operatorname{tr}\left[{(\gamma^{\beta}_{A})}^{-1}\right].$ (14)
If the Hamiltonian of $A$ is trivial or in general $\widehat{H}_{A}=c\mathbbm{1}_{A}$ for $c\in\mathbbm{R}$, then $F^{\beta}_{\infty}[\mathcal{U}]=2\beta^{-1}\ln|A|$.
The proof is given in Appendix A.6.
Lemma 1.
The sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]$, $\alpha\in[\frac{1}{2},\infty)$, is nonincreasing under the action of Gibbs-subpreserving superchannel. It remains invariant under the action of Gibbs preserving unitary superchannels.
The proof is given in Appendix A.2.
Lemma 2.
Let $\mathcal{N}$ be a quantum channel. The generalized free energies obey order $\mathbf{F}^{\beta,(x)}[\mathcal{N}]\leq\mathbf{F}^{\beta,(y)}[\mathcal{N}]$ if respective generalized state divergence obey the same order, i.e., $\mathbf{D}^{(x)}(\rho\|\sigma)\leq\mathbf{D}^{(y)}(\rho\|\sigma)$ for quantum states $\rho,\sigma$.
Proof.
The proof follows from the definition of the generalized free energy and proof argument similar to [30, Lemma 1]. ∎
A direct consequence of the above lemma is that, for all $\alpha_{1},\alpha_{2}\in[\frac{1}{2},\infty)$ such that $\alpha_{1}\leq\alpha_{2}$, we have $F^{\beta}_{\alpha_{1}}[\mathcal{N}]\leq F^{\beta}_{\alpha_{2}}[\mathcal{N}]$.
Lemma 3.
The free energy $F^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is lower bounded by the completely-bounded norm of the channel with the absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$,
$F^{\beta}[\mathcal{N}]\geq\frac{1}{2\beta}\norm{\mathcal{N}-\mathcal{T}^{\beta}}_{\diamond}^{2}.$ (15)
Proof.
The proof follows from Pinsker’s inequality for any two states $\rho,\sigma$:
$D(\rho\|\sigma)\geq\frac{1}{2}\norm{\rho-\sigma}_{1}^{2}.$ (16)
The Pinsker’s inequality implies that, for any two quantum channels $\mathcal{N},\mathcal{M}\in\operatorname{Ch}(A^{\prime},A)$,
$D[\mathcal{N}\|\mathcal{M}]\geq\frac{1}{2}\norm{\mathcal{N}-\mathcal{M}}_{\diamond}^{2}.$ (17)
IV Dynamical Resource Theory of Athermality
We now introduce framework and characterization of the thermodynamic resource theory of square quantum channels. A quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is a square quantum channel if $|A^{\prime}|=|A|$; in other words, a quantum channel with square Choi operator is called a square channel. To quantify the thermodynamic resourcefulness of a quantum channel $\mathcal{N}$, we consider its distinguishability from a thermal channel $\mathcal{T}^{\beta}$ whose output $\gamma^{\beta}$ is in thermal equilibrium with the bath [31]. The absolutely thermal channel $\mathcal{T}^{\beta}$ represents thermalization process with respect to the given bath at inverse temperature $\beta$.
To characterize a resource theory of athermality of quantum channels, we identify our free object to be the absolutely thermal channel $\mathcal{T}^{\beta}$ and the free operations to be Gibbs preserving superchannels (GPSC). This is a relevant assumption as thermalization processes seem to be natural in practical scenarios where quantum systems inherently tend to thermalize upon interaction with the bath. The closed systems in isolation with the bath are nontrivial and idealistic. We write the athermal channel resource as $(\mathcal{N},\mathcal{T}^{\beta})$ to specify that the athermality of the channel $\mathcal{N}$ is being considered with respect to $\mathcal{T}^{\beta}$. We call two channel-resources $(\mathcal{N},\mathcal{T}^{\beta})$ and $(\mathcal{M},\mathcal{T}^{\beta^{\prime}})$ to be equivalent, $(\mathcal{N},\mathcal{T}^{\beta})\sim(\mathcal{M},\mathcal{T}^{\beta^{\prime}})$, if they are interconvertible through some GPSCs.
For a quantum superchannel $\Theta\in\operatorname{SCh}((A^{\prime},A),(B^{\prime},B))$ with the following action,
$\Theta(\mathcal{N}_{A^{\prime}\to A})=\mathcal{Q}_{A\to B}\circ\mathcal{N}\circ\mathcal{P}_{B^{\prime}\to A^{\prime}},$ (18)
$\Theta$ is Gibbs preserving if and only if $\mathcal{Q}_{A\to B}$ is Gibbs preserving channel, i.e., $\mathcal{Q}(\gamma^{\beta}_{A})=\gamma^{\beta}_{B}$. A quantum superchannel $\widetilde{\Theta}\in\operatorname{SCh}((A^{\prime},A),(B^{\prime},B))$, where $\widetilde{\Theta}(\mathcal{N})=\widetilde{\mathcal{Q}}_{PA\to B}\circ\mathcal{N}\circ\widetilde{\mathcal{P}}_{B^{\prime}\to PA^{\prime}}$ is also Gibbs preserving if its postprocessing channel $\widetilde{\mathcal{Q}}_{PA\to B}$ is absolutely thermal.
Golden units of athermality.— For a square quantum channel $\mathcal{N}_{A^{\prime}\to A}$, the free energy is maximum if and only if the channel is unitary. For an arbitrary pair $(\mathcal{U}_{A^{\prime}\to A},\mathcal{V}_{A^{\prime}\to A})$ of unitary channels, there always exists a GPSC $\Theta^{\beta}\in\operatorname{SCh}((A^{\prime},A),(A^{\prime},A))$ such that $\Theta^{\beta}(\mathcal{U})=\mathcal{V}$, for an instance, $\Theta^{\beta}(\mathcal{N})=\mathcal{N}\circ(\mathcal{U}^{{\dagger}}\circ\mathcal{V})$. That is, all unitary channels $(\mathcal{U},\mathcal{T}^{\beta})$ are equally resourceful, $(\operatorname{id},\mathcal{T}^{\beta})\sim(\mathcal{U},\mathcal{T}^{\beta})$. This is consistent with the free energy of all unitary channels $(\mathcal{U},\mathcal{T}^{\beta})$ being the same. The free energy of a unitary channel is least when the output Hamiltonian is fully degenerate (see Lemma 9). In other words, among all unitary resource channels $(\mathcal{U}_{m},\mathcal{T}^{\beta})$ for a fixed $m$, the unitary resource channels $(\mathcal{U}_{m},\mathcal{R}^{\pi})$ have minimum free energy. With freedom of choice for golden units and the fact that GPSCs are free operations in our athermality framework, we mark our golden unit as the identity resource channel $(\operatorname{id}_{m},\mathcal{R}^{\pi})$, where $\mathcal{R}^{\pi}_{A^{\prime}\to A}$ is the absolutely thermal channel for fully degenerate output Hamiltonian ($\widehat{H}_{A}\propto\mathbbm{1}_{A}$).
It is evident that $(\operatorname{id}_{A^{\prime}\to A},\mathcal{R}^{\pi}_{A^{\prime}\to A})$ is more resourceful than $(\operatorname{id}_{B^{\prime}\to B},\mathcal{R}^{\pi}_{B^{\prime}\to B})$ if $|A|>|B|$ [24, 31, 25], therefore, we will have to be considerate of the dimension of $\operatorname{id}$ in the golden unit. For a golden unit $(\operatorname{id}_{m},\mathcal{R}^{\pi})$, where the subscript index $m$ is to denote that $\operatorname{id}$ acts on $m$-dimensional Hilbert space. The identity channel $\operatorname{id}_{m}$ (or any equivalent unitary channel $\mathcal{U}_{m}$) allows for the perfect transmission of $m^{2}$-dimensional quantum states [58, 59], equivalent to the distribution of a maximally entangled state $\Phi$ of Schmidt rank $m$ [60, 59, 27].
The uniformly mixing channel $\mathcal{R}^{\pi}_{A^{\prime}\to A}$, $|A^{\prime}|=|A|=m$, can be obtained by the action of unitary channels $\mathcal{U}(\cdot)=U(\cdot){U}^{\dagger}$ uniformly picked at random with respect to Haar measure,
$\mathcal{R}^{\pi}(\cdot)=\int_{U\in\mathbbm{U}(m)}{\operatorname{d}\!\mu_{U}}U(\cdot)U^{\dagger},$ (19)
where $\mathbbm{U}(m)$ is the set of all unitary operators $U_{A^{\prime}\to A}$. We can also obtain $\mathcal{R}^{\pi}$ by taking a uniform mixture of Weyl unitary channels $\mathcal{W}^{i}(\cdot)=W^{i}(\cdot){W^{i}}^{\dagger}$,
$\mathcal{R}^{\pi}(\cdot)$ $=\frac{1}{m^{2}}\sum_{i=0}^{m^{2}-1}{W}^{i}(\cdot){W^{i}}^{\dagger},$ (20)
where $\{W^{i}\}_{i=0}^{m^{2}-1}$ is the set (group) of Weyl unitaries; $\{W^{i}\}_{i=0}^{m^{2}-1}$ forms a complete orthonormal basis for the space of all linear operators acting on $m$-dimensional Hilbert space. Let $\mathcal{W}^{0}=\mathrm{\operatorname{id}}_{m}$ and $\operatorname{id}_{m}^{\perp}:=\sum_{i=1}^{m^{2}-1}\mathcal{W}^{i}$, then
$\mathcal{R}^{\pi}$ $=\frac{1}{m^{2}}\operatorname{id}_{m}+\frac{1}{m^{2}}\sum_{i=1}^{m^{2}-1}\mathcal{W}^{i}=\frac{1}{m^{2}}(\operatorname{id}_{m}+\operatorname{id}_{m}^{\perp}).$ (21)
This implies (cf. [61, Proposition 11])
$\frac{1}{2}\norm{\operatorname{id}_{m}-\mathcal{R}^{\pi}}_{\diamond}=1-\frac{1}{m^{2}}.$ (22)
Channel conversion distance.– For a given resource channel $(\mathcal{N},\mathcal{T}^{\beta_{1}})$ being transformed to another resource channel $(\mathcal{M},\mathcal{T}^{\beta_{2}})$, the conversion distance under GPSCs are defined as
$d_{\mathrm{GP}}((\mathcal{N},\mathcal{T}^{\beta_{1}})\rightarrow(\mathcal{M},\mathcal{T}^{\beta_{2}}))$
$\qquad=\min_{\Theta}\left\{\frac{1}{2}\|\mathcal{M}-\Theta(\mathcal{N})\|_{\diamond}\penalty 10000\ :\penalty 10000\ \Theta(\mathcal{T}^{\beta_{1}})=\mathcal{T}^{\beta_{2}}\right\}.$ (23)
Definition 2 (One-shot athermality distillation and formation).
For an error $\varepsilon\in[0,1]$, the single-shot athermality distillation of the resource channel $(\mathcal{N},R^{\gamma})$ under GPSCs is defined as
[…]
$:=\sup_{m}\left\{\ln m:\penalty 10000\ d_{\mathrm{GP}}((\mathcal{N},\mathcal{T}^{\beta})\rightarrow(\operatorname{id}_{m},\mathcal{R}^{\pi}))\leq\varepsilon\right\}.$ (24)
For an error $\varepsilon\in[0,1]$, the single-shot athermality formation (cost) of the resource channel $(\mathcal{N},\mathcal{R}^{\gamma})$ is defined as
[…]
$:=\inf_{m}\left\{\ln m:\penalty 10000\ d_{\mathrm{GP}}((\operatorname{id}_{m},\mathcal{R}^{\pi})\rightarrow(\mathcal{N},\mathcal{T}^{\beta}))\leq\varepsilon\right\}.$ (25)
The single-shot athermality distillation $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ provides the largest dimensional golden unit $(\operatorname{id}_{m},\mathcal{R}^{\pi})$ that can be distilled from the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs, up to allowed error $\varepsilon$. We could possibly also interpret $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ as the maximum number of elementary golden units $(\operatorname{id}_{2},\mathcal{R}^{\pi})$ distillable from the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs. The single-shot athermality formation $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ provides the least dimensional golden unit $(\operatorname{id}_{m},\mathcal{R}^{\pi})$ needed to form the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs, up to allowed error $\varepsilon$. We could possibly also interpret $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ as the minimum number of elementary golden units $(\operatorname{id}_{2},\mathcal{R}^{\pi})$ spent to form the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs.
Theorem 3 (Athermality distillation and formation).
For any error $\varepsilon\in[0,1]$ and a given resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, the single-shot athermality distillation and formation are proportional to the $\varepsilon$-hypothesis-testing free energy and the $\varepsilon$-max-free energy of the channel $\mathcal{N}$, respectively,
$\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ $=\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (26)
$\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ $=\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (27)
$F^{\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
We provide the detailed proof of the theorem in Appendix A.4. The zero-error cost is proportional to the channel max-free energy $F_{\infty}^{\beta}[\mathcal{N}]$, that can be expressed as a semidefinite program, see Appendix A.5.
The one-shot athermality distillation and formation of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ are associated with how well the channel $\mathcal{N}$ can be discriminated from $\mathcal{T}^{\beta}$ in an asymmetric channel discrimination task [62, 63, 64, 65], where $\mathcal{N}$ is null hypothesis and $\mathcal{T}^{\beta}$ is alternate hypothesis. Broadly speaking, there two different strategies one can apply when multiple uses of channels are allowed, namely, adaptive and nonadaptive [62, 65]. Channel discrimination under nonadaptive strategy can be thought of as a special case of discrimination task with adaptive strategy. Nonadaptive strategy allows for parallel uses of a channel with free preprocessing and postprocessing operations, while adaptive strategy allows for sequential uses of channel with successive uses of channel being interleaved with a free adaptive channel [62, 63].
It intuitively holds that the asymptotic resource distillation rate is lesser than or equal to the formation rates in a resource theory where golden units (most resourceful objects) are distilled from lesser resource objects and spent as costs to form lesser resource objects (cf. [66]). It follows from the relation between the min relative entropy and max relative entropy, $D^{0}_{\rm H}(\rho\|\sigma)\leq D_{\infty}(\rho\|\sigma)$, that the zero-error athermality distillation and formation satisfy $\mathrm{Dist}^{0}(\mathcal{N},\mathcal{T}^{\beta})\leq\mathrm{Cost}^{0}(\mathcal{N},\mathcal{T}^{\beta})$. We now provide the trade-off relation between the single-shot athermality distillation and formation in the following proposition.
Proposition 4 (Thermodynamic yield-cost trade-off relation).
For a quantum channel $\mathcal{N}$ and $\varepsilon\in(0,1)$, the distillation and formation of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ satisfy the following trade-off relation,
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{1-\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right).$ (28)
The proof is given in Appendix A.7.
Based on the assistance of adaptive and nonadaptive strategies for the channel discrimination tasks (see [62, 65] for formal discussion), for an error $\varepsilon\in(0,1)$, we define the adaptive ($\rm ad$}) and nonadaptive ($\parallel$) asymptotic athermality distillation and formation rates of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ as
$\mathscr{C}_{\mathrm{distill}}^{{\varepsilon,\rm ad}}[\mathcal{N}]$ $:=\limsup_{n\to\infty}\frac{1}{n}\,\mathrm{Distill}^{\varepsilon}_{\rm ad}\left(\mathcal{N}^{\otimes n},\left(\mathcal{T}^{\beta}\right)^{\otimes n}\right),$ (29)
$\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$ $:=\liminf_{n\to\infty}\frac{1}{n}\,\mathrm{Cost}^{\varepsilon}_{\rm ad}\left(\mathcal{N}^{\otimes n},\left(\mathcal{T}^{\beta}\right)^{\otimes n}\right),$ (30)
$\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]$ $:=\limsup_{n\to\infty}\frac{1}{n}\,\mathrm{Distill}^{\varepsilon}_{\parallel}\left(\mathcal{N}^{\otimes n},\left(\mathcal{T}^{\beta}\right)^{\otimes n}\right),$ (31)
$\mathscr{C}^{\varepsilon,\parallel}_{\mathrm{cost}}[\mathcal{N}]$ $:=\liminf_{n\to\infty}\frac{1}{n}\,\mathrm{Cost}^{\varepsilon}_{\parallel}\left(\mathcal{N}^{\otimes n},\left(\mathcal{T}^{\beta}\right)^{\otimes n}\right).$ (32)
It directly follows from the definitions that $\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]\leq\mathscr{C}_{\mathrm{distill}}^{\varepsilon,{\rm ad}}[\mathcal{N}]$ and $\mathscr{C}^{\varepsilon,\parallel}_{\mathrm{cost}}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$, and $\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,\parallel}_{\mathrm{cost}}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$. We prove that the asymptotic distillation and formation rates are equal.
Theorem 4 (Asymptotic reversibility).
The resource theory of athermality of quantum channels with GPSCs as free operations is asymptotically reversible, for any $\varepsilon\in(0,1)$,
$\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]$ $=\mathscr{C}_{\mathrm{cost}}^{\varepsilon,\parallel}[\mathcal{N}]=\mathscr{C}_{\mathrm{distill}}^{\varepsilon,{\rm ad}}[\mathcal{N}]=\mathscr{C}_{\mathrm{cost}}^{\varepsilon,{\rm ad}}[\mathcal{N}]$
$=\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]=\frac{\beta}{2}F^{\beta}[\mathcal{N}].$ (33)
Proof.
For any $\varepsilon\in(0,1)$, we have from [62, Theorem 1],
$\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]=\mathscr{C}_{\mathrm{distill}}^{\varepsilon,{\rm ad}}[\mathcal{N}]=\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}],$ (34)
and from [57, Theorem 4.1],
$\mathscr{C}_{\mathrm{cost}}^{\varepsilon,\parallel}[\mathcal{N}]=\mathscr{C}_{\mathrm{cost}}^{\varepsilon,{\rm ad}}[\mathcal{N}]=\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}].$ (35)
We conclude the proof by recalling that $F^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$. ∎
Both the athermality distillation and formation rates of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ converge to the same value, which is half the relative entropy between the channel and the absolutely thermal channel $\mathcal{T}^{\beta}$. Therefore, the resource theory of athermality under GPSCs is reversible. The asymptotic athermality distillation and formation rates of a tensor-product quantum resource channel $(\mathcal{N}\otimes\mathcal{M},\mathcal{T}^{\beta})$, where the channel outputs of quantum channels $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{B^{\prime}\to B}$ are noninteracting and $\mathcal{T}^{\beta}_{A^{\prime}B^{\prime}\to AB}=\mathcal{T}^{\beta}_{A^{\prime}\to A}\otimes\mathcal{T}^{\beta}_{B^{\prime}\to B}$, are equal to $\frac{\beta}{2}\left(F^{\beta}[\mathcal{N}]+F^{\beta}[\mathcal{M}]\right)$. This is a straightforward implication of the additivity of the free energy under tensor-product channels with noninteracting channel outputs.
IV.1 Remarks on the resource theory of athermality
We note that in [19], the restricted form of the free energy associated to athermal channel, where reference to the channel are not considered, is shown to be related to the maximal work distillation. In particular, they showed that for the asymptotically many uses of each channel, the distillable work is an additive function of the considered channels. The allowed free operations for their task are thermal operations, a strict subset of Gibbs preserving channels. Our approach here can be thought of as a generalization where access to a reference to the channel is allowed along with adaptive strategies for the uses of the channel in the distillation (and formation) tasks. The access to the channel reference and adaptive strategy for distillation (and formation) tasks provide a general approach in resource-theoretic framework that involves channel discrimination. The rates of distillation and formation will never be less than that of the restricted case with parallel uses of channels and no allowance of the reference to the channel. The sandwiched Rényi free energy is additive for tensor-product channels with noninteracting outputs for $\alpha\in[\frac{1}{2},\infty)$, including the free energy ($\alpha=1$). This additivity property of the free energy makes the asymptotic rates of distillation and formation additive in our resource theory of athermality.
[…]
The dynamical thermodynamic resource theory exhibit some features beyond the the static thermodynamic resource theory. When Hamiltonian $\widehat{H}_{A}$ is non-degenarate, i.e., no two distinct energy eigenstates have same eigenvalues, then a state has the maximum free energy if and only if it is the most excited energy eigenstate (eigenstate with the highest energy eigenvalue). The state with maximum free energy for a quantum system with non-degenerate Hamiltonian is unique. This is characteristically different from the dynamical case where all the unitary channels $\mathcal{U}_{A^{\prime}\to A}$ have maximum free energy. We do have the resource theory of entanglement of states where there are many golden units called maximally entangled states [67], which are all equivalent to each other up to the action of local unitary channels.
The conversion of a resource channel $(\mathcal{N}_{A^{\prime}\to A},\mathcal{T}^{\beta}_{A^{\prime}\to A})$ to $(\mathcal{T}^{\beta}_{A^{\prime}\to A},\mathcal{T}^{\beta}_{A^{\prime}\to A})$ is an informational thermodynamic task of erasing the channel $\mathcal{N}$ via thermalization [68, 69, 38, 70], whereas the conversion of the absolutely thermal channel $(\mathcal{T}^{\beta},\mathcal{T}^{\beta})$ to $(\mathcal{N},\mathcal{T}^{\beta})$ is an informational thermodynamic task of preparing or forming the channel.
V Informational aspects of free energy
We discuss connection of the dynamical free energy with the channel capacity and the free energy of the channel output. We observe connections of the resource theory of athermality with the information processing tasks of private randomness and purity distillation.
The quantum mutual information $I(A;B)_{\rho}$ of a quantum state $\rho_{AB}$ is defined as
$I(A;B)_{\rho}$ $=D(\rho_{AB}\|\rho_{A}\otimes\rho_{B})$
$=S(A)_{\rho}+S(B)_{\rho}-S(AB)_{\rho}.$ (36)
It is the minimal rate of noise needed to erase the total correlations in the state $\rho_{AB}$ so that the result is a product state $\rho_{A}\otimes\rho_{B}$ [71].
The quantum mutual information $I(\mathcal{N},\rho)$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ and an input state $\rho_{RA^{\prime}}$ to $\operatorname{id}_{R}\otimes\mathcal{N}$ is defined as the quantum mutual information $I(R;A)_{\mathcal{N}(\rho_{RA^{\prime}})}$, $I(\mathcal{N},\rho):=I(R,A)_{\mathcal{N}(\rho_{RA^{\prime}})}=D(\mathcal{N}(\rho_{RA^{\prime}})\|\rho_{R}\otimes\mathcal{N}(\rho_{A^{\prime}}))$. The quantum mutual information of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as
$I[\mathcal{N}]:=\sup_{\psi\in\operatorname{St}(RA^{\prime})}I(R;A)_{\mathcal{N}(\psi_{RA^{\prime}})},$ (37)
where it suffices to optimize over pure state $\psi_{RA^{\prime}}$. The quantum mutual information of a channel indicates how well the channel can preserve the total correlation (quantum+classical) between the reference and the input. It finds operational meaning as the entanglement-assisted classical capacity of the channel [72] (see also [26, 63] for connection with quantum rebound capacity) and the asymptotic quantum simulation cost under no-signalling assisted codes [73].
Proposition 5.
The free energy $F^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ satisfies
$F^{\beta}[\mathcal{N}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[\beta^{-1}I(R;A)_{\mathcal{N}(\psi)}+F^{\beta}(\mathcal{N}(\psi_{A^{\prime}}))\right],$ (38)
where it suffices to optimize over pure states $\psi_{RA^{\prime}}$, $|R|=|A^{\prime}|$.
Proof.
The proof follows from the following observations. $F[\mathcal{N}]=\beta^{-1}\sup_{\psi\in\operatorname{St}(RA^{\prime})}D(\mathcal{N}(\psi_{RA^{\prime}})\|\psi_{R}\otimes\gamma^{\beta}_{A})$, where it suffices to optimize over pure states $\psi_{RA^{\prime}}$, and
$\beta^{-1}\sup_{\psi\in\operatorname{St}}D(\mathcal{N}(\psi_{RA^{\prime}})\|\psi_{R}\otimes\gamma^{\beta}_{A})$
$=$ $\beta^{-1}\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[D(\psi^{\mathcal{N}}_{RA}\|\psi_{R}\otimes\psi^{\mathcal{N}}_{A})+D(\psi^{\mathcal{N}}_{A}\|\gamma^{\beta}_{A})\right]$
$=$ $\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[\beta^{-1}I(R;A)_{\mathcal{N}(\psi_{RA^{\prime}})}+F(\mathcal{N}(\psi_{A^{\prime}}))\right],$ (39)
where $\psi^{\mathcal{N}}_{RA}:=\mathcal{N}(\psi_{RA^{\prime}})$. ∎
The following corollary is a direct from the above proposition.
Corollary 1.
The free energy ${F}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is lower bounded as
$F^{\beta}[\mathcal{N}]\geq\beta^{-1}I(R;A)_{\Phi^{\mathcal{N}}}+F(\mathcal{N}(\pi_{A^{\prime}})),$ (40)
and upper bounded as
$F^{\beta}[\mathcal{N}]\leq\beta^{-1}I[\mathcal{N}]+\sup_{\psi\in\operatorname{St}(A^{\prime})}F^{\beta}(\mathcal{N}(\psi)).$ (41)
The free energy of a unitary channel $\mathcal{U}_{A^{\prime}\to A}$ is lower bounded as
$F^{\beta}[\mathcal{U}]\geq\beta^{-1}\left(\ln|A|+\ln Z^{\beta}_{A}\right)+\frac{1}{|A|}\operatorname{tr}[\widehat{H}_{A}],$ (42)
where $Z^{\beta}_{A}:=\operatorname{tr}[\exp(-\beta\widehat{H}_{A})]$.
Proposition 6.
Given a unitary channel $\mathcal{U}_{A^{\prime}\to A}$, the minimum generalized free energy of the channel is achieved for the output system Hamiltonian $\widehat{H}_{A}=c\mathbbm{1}_{A}$, In particular, for $\alpha\in[1,\infty)$,
$\min_{\widehat{H}_{A}}F^{\beta}_{\alpha}[\mathcal{U}]=2\beta^{-1}\ln|A|,\penalty 10000\ \forall\penalty 10000\ \alpha\in[1,\infty).$ (43)
Proof.
From Corollary 1, we have $F^{\beta}[\mathcal{U}]\geq f(\widehat{H}_{A})$ where
$f(\widehat{H}_{A})=\beta^{-1}(\ln|A|+\ln Z_{A}^{\beta}+\frac{\beta}{|A|}\operatorname{tr}(\widehat{H}_{A})).$ (44)
To find the global minima of the above operator function, we analyze the operator derivative $\nabla_{\widehat{H}_{A}}f(\widehat{H}_{A})$. Using arguments similar to Lemma 9 in Appendix A.8, we can show that
$\min_{\widehat{H}_{A}}f(\widehat{H}_{A})=f(c\mathbbm{1}_{A})=2\beta^{-1}\ln|A|.$ (45)
Combining the above result with that of Lemma 9 and Lemma 2, we have the following for $\alpha\in(1,\infty)$,
$2\beta^{-1}\ln|A|\leq$ $\min_{\widehat{H}_{A}}F^{\beta}[\mathcal{U}]\leq\min_{\widehat{H}_{A}}F_{\alpha}^{\beta}[\mathcal{U}]$ (46)
$\leq$ $\min_{\widehat{H}_{A}}F_{\infty}^{\beta}[\mathcal{U}]=2\beta^{-1}\ln|A|$
$\implies\min_{\widehat{H}_{A}}F^{\beta}[\mathcal{U}]=$ $\min_{\widehat{H}_{A}}F_{\alpha}^{\beta}[\mathcal{U}]=\min_{\widehat{H}_{A}}F_{\infty}^{\beta}[\mathcal{U}]=2\beta^{-1}\ln|A|.$ (47)
Private randomness distillation.— The private randomness capacity of a channel $\mathcal{N}_{A^{\prime}\to A}$ is the maximum rate, in an asymptotic setting of i.i.d. uses of the channel, at which the receiver accessing output $A$ can extract private randomness, against an eavesdropper accessing extension $E$ of an isometric channel extension $\mathcal{V}^{\mathcal{N}}_{A^{\prime}\to AE}$ of the channel N, when the sender sends states through the channel [35]. The private randomness capacity $P[\mathcal{N}]=D[\mathcal{N}\|\mathcal{R}^{\pi}]=\lim_{\beta\to 0^{+}}D[\mathcal{N}\|\mathcal{T}^{\beta}]$. The private randomness capacity of a square quantum channel is twice its asymptotic thermodynamic distillation or formation rate under the action of GPSCs in the limit inverse temperature tends to vanish.
Dynamical resource theory of purity.— By considering $\beta\to 0^{+}$ or the Hamiltonian of the channel output to be proportional to $\mathbbm{1}$ in our dynamical resource theory of athermality, we obtain the dynamical resource theory of purity [35, 74, 75, 36] where unitary channels are golden units and the uniformly mixing channels are free objects. When $\widehat{H}_{A}\propto\mathbbm{1}_{A}$, then $\mathcal{T}^{\beta}_{A^{\prime}\to A}=\mathcal{R}^{\pi}_{A^{\prime}\to A}$ as $\gamma^{\beta}_{A}=\pi_{A}$ due to $\widehat{H}_{A}$ being fully degenerate Hamiltonian, i.e., all energy eigenstates have the same eigenvalue. The asymptotic purity distillation and formation rate of a square quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is equal to $\frac{1}{2}D[\mathcal{N}\|\mathcal{R}^{\pi}]=\frac{1}{2}\lim_{\beta\to 0^{+}}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ [35, 24, 36].
Entropy and thermal entropy.— The entropy $S[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ and its private randomness capacity $P[\mathcal{N}]$ satisfy the following duality or trade-off relation [30, 31]:
$S[\mathcal{N}]+P[\mathcal{N}]=\ln|A|.$ (48)
We have a similar dual relationship between the thermal entropy $S^{\beta}[\mathcal{N}]:=\ln Z^{\beta}_{A}-D[\mathcal{N}\|\mathcal{T}^{\beta}]$ [25] and the resource-theoretic free energy $F^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$:
$S^{\beta}[\mathcal{N}]+\beta F^{\beta}[\mathcal{N}]=\ln Z^{\beta}_{A},$ (49)
where $Z^{\beta}_{A}=\operatorname{tr}[\exp(-\beta\widehat{H}_{A})]$. For the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{R}^{\pi}]$, where $|A|<\infty$ and $\alpha\in[\frac{1}{2},\infty)$, we have
$\lim_{\beta\to 0^{+}}D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]=D_{\alpha}[\mathcal{N}\|\mathcal{R}^{\pi}]=\ln|A|-S_{\alpha}[\mathcal{N}].$ (50)
If we consider the Hamiltonian of the output $A$ of the channel $\mathcal{N}_{A^{\prime}\to A}$ to be trivial, $\widehat{H}_{A}=0$ and $\widehat{H}_{RA}=\widehat{H}_{R}\otimes\mathbbm{1}_{A}$, then for any $\beta>0$ and $\forall\penalty 10000\ \alpha\in[\frac{1}{2},\infty)$,
$S_{\alpha}[\mathcal{N}]+\beta F^{\beta}_{\alpha}[\mathcal{N}]=\ln|A|.$ (51)
The min-entropy $S_{\infty}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is associated with the decoupling ability of the channel [38, Theorem 1 & Proposition 1] and the environment-assisted erasure cost $W[A|E]_{\mathcal{N}}$ of the output $A$ of the channel when access to the environment $E$ is provided to the erasure [38, Theorem 3 & Proposition 2]. These operational interpretations of the min-entropy of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ directly provides operational interpretations for given information processing tasks as the max-free energy of the channel $F^{\beta}_{\infty}[\mathcal{N}]=\beta^{-1}(\ln|A|-S_{\infty}[\mathcal{N}])$ for $\widehat{H}_{A}=0$.
VI Thermodynamical free energy and work extraction
For a quantum state $\rho_{A}$, the entropy $S(\rho)$ is $S(\rho):=S(A)_{\rho}:=-D(\rho_{A}\|\mathbbm{1}_{A})$ and the energy $E(\rho)$ is $E(\rho)=\langle\widehat{H}_{A}\rangle_{\rho}=\operatorname{tr}[\widehat{H}_{A}\rho_{A}]$. For a fixed inverse temperature $\beta$, the thermodynamic non-equilibrium free energy $F^{\beta}_{c}(\rho)$ of a quantum state $\rho_{A}$ is related to its entropy and energy as
$F^{\beta}_{c}(\rho)=E(\rho)-\beta^{-1}S(\rho)=\beta^{-1}D(\rho\|\widehat{\gamma}^{\beta}),$ (52)
where $\widehat{\gamma}^{\beta}_{A}=\exp(-\beta\widehat{H}_{A})$ is the thermal operator (unnormalized thermal state).
The entropy $S[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as [56, 76, 24, 25]
$S[\mathcal{N}]$ $:=-D[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]$
$=\inf_{\psi\in\operatorname{St}(RA^{\prime})}[S(RA)_{\mathcal{N}(\psi)}-S(R)_{\psi}],$ (53)
where it suffices to optimize only over pure states $\psi_{RA^{\prime}}$.
Definition 3 (Thermodynamic free energy).
The generalized thermodynamic free energy $\mathbf{F}^{\beta}_{c}[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as
$\mathbf{F}^{\beta}_{c}[\mathcal{N}]:=\beta^{-1}\mathbf{D}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}],$ (54)
where $\widehat{\mathcal{T}}^{\beta}=\mathcal{R}^{\widehat{\gamma}^{\beta}}$, i.e., $\widehat{\mathcal{T}}^{\beta}(\rho_{A^{\prime}})=\operatorname{tr}[\rho_{A^{\prime}}]\widehat{\gamma}^{\beta}_{A}$ is a replacer map that is completely positive but not trace-preserving. The thermodynamic non-equilibrium free energy $F^{\beta}_{c}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is $F^{\beta}_{c}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$, obtained by taking generalized channel divergence to be relative entropy.
The function $\mathbf{D}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]=-\mathbf{S}^{\beta}[\mathcal{N}]$, where $\mathbf{S}^{\beta}[\mathcal{N}]$ is called the generalized thermal entropy of the channel $\mathcal{N}$ [25]. We see that $\mathcal{F}^{\beta}_{c}[\mathcal{N}]=\beta^{-1}\mathcal{S}^{\beta}[\mathcal{N}]$. The thermodynamic sandwiched Rényi free divergence $F^{\beta}_{c,\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\widehat{T}^{\beta}]$ for $\alpha\in[\frac{1}{2},1)\cup(1,\infty]$. The sandwiched Rényi free energy of the absolutely thermal channel reduces to the Helmholtz free energy,
$F^{\beta}_{c,\alpha}[\widehat{T}^{\beta}]=-\ln Z^{\beta}_{A},$ (55)
for $\alpha\in[\frac{1}{2},\infty)$, where $\alpha=1$ is taken as limit $\alpha\to 1$. For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ with $\widehat{H}_{A}\propto\mathbbm{1}_{A}$ and noninteracting reference to the channel output, the thermodynamic sandwiched Rényi free energy reduces to the sandiwched Rényi entropy (up to a multiplicative factor of the negative inverse temperature),
$F_{c,\alpha}[\mathcal{N}]=-\beta^{-1}S_{\alpha}[\mathcal{N}].$ (56)
The resource-theoretic sandwiched Rényi free energy and the thermodynamic sandwiched Rényi free energy of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ are related as
$F^{\beta}_{\alpha}[\mathcal{N}]=F^{\beta}_{c,\alpha}[\mathcal{N}]+\beta^{-1}\ln Z^{\beta}_{A},$ (57)
for $\alpha\in[\frac{1}{2},\infty]$. $F^{\beta}_{c}[\mathcal{N}]$ satisfies all the axiomatic properties (A1-A6) as desired from the free energy function. We also have
$F^{\beta}[\mathcal{N}]$ $=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$
$=\sup_{\psi\in\operatorname{St}(RA^{\prime})}[F^{\beta}(\mathcal{N}(\psi_{RA^{\prime}}))-F^{\beta}(\psi_{R})],$ (58)
where it suffices to optimize only over pure states $\psi_{RA^{\prime}}$. The thermodynamic sandwiched Rényi free divergence $F_{c,\alpha}[\mathcal{N}]$, for $\alpha\in[\frac{1}{2},1)\cup(1,\infty)$ satisfies the axiomatic properties (A1A5)(A1-A5) and quasi-convexity.
The following lemma follows from the definitions of the resource-theoretic and thermodynamic free energies.
Lemma 4.
The resource-theoretic sandwiched Rényi free energy $F_{\alpha}[\mathcal{N}]$, for $\alpha\in[\frac{1}{2},\infty)$, is equal to the difference of the thermodynamic sandwiched Rényi free energy of the channel and the absolutely thermal channel,
$F^{\beta}_{c,\alpha}[\mathcal{N}]-F^{\beta}_{c,\alpha}[\mathcal{T}^{\beta}]=F^{\beta}_{\alpha}[\mathcal{N}].$ (59)
Based on the “channelized” entropy and free energy definitions, we define an energy function called the thermodynamic energy $E[\mathcal{N}]$ of a quantum channel $\mathcal{N}$.
Definition 4.
The thermodynamic energy $E[\mathcal{N}]$ of a quantum channel $\mathcal{N}$ as
$E[\mathcal{N}]:=\sup_{\psi\in\operatorname{St}(RA^{\prime})}[E(\mathcal{N}(\psi_{RA^{\prime}}))-E(\psi_{R})],$ (60)
where it suffices to optimize only over pure states $\psi_{RA^{\prime}}$ and $|R|=|A^{\prime}|$.
Proposition 7.
The thermodynamic energy $E[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, with the reference-output Hamiltonian as $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes\widehat{H}_{A}+\widehat{H}_{R}\otimes\mathbbm{1}_{A}+\widehat{H}^{\rm int}_{RA}$, is
$E[\mathcal{N}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[\langle\widehat{H}_{A}\rangle_{\mathcal{N}(\psi_{A^{\prime}})}+\langle\widehat{H}^{\rm int}_{RA}\rangle_{\mathcal{N}(\psi_{RA^{\prime}})}\right],$ (61)
where it suffices to optimize over pure states $\psi_{RA^{\prime}}$.
Proof.
The proof follows from the definition of $E[\mathcal{N}]$ and the fact that, for state $\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})$ and Hamiltonian $\widehat{H}_{RA}$, we have
$E(\mathcal{N}(\psi_{RA^{\prime}}))-E(\psi_{R})$
$=$ $\operatorname{tr}[\widehat{H}_{RA}\mathcal{N}(\psi_{RA^{\prime}})]-\operatorname{tr}[\widehat{H}_{R}\psi_{R}]$
$=$ $\operatorname{tr}[(\widehat{H}_{R}\otimes\mathbbm{1}_{A})\mathcal{N}(\psi_{RA^{\prime}})]+\operatorname{tr}[(\mathbbm{1}_{R}\otimes\widehat{H}_{A})\mathcal{N}(\psi_{RA^{\prime}})]$
$+\operatorname{tr}[\widehat{H}_{RA}^{\rm int}\mathcal{N}(\psi_{RA^{\prime}})]-\operatorname{tr}[\widehat{H}_{R}\psi_{R}]$
$=$ $\operatorname{tr}[\widehat{H}_{R}\psi_{R}]+\operatorname{tr}[\widehat{H}_{A}\mathcal{N}(\psi_{A^{\prime}})]+\operatorname{tr}[\widehat{H}^{\rm int}_{RA}\mathcal{N}(\psi_{RA^{\prime}})]$
$-\operatorname{tr}[\widehat{H}_{R}\psi_{R}]$
$=$ $\operatorname{tr}[\widehat{H}_{A}\mathcal{N}(\psi_{A^{\prime}})]+\operatorname{tr}[\widehat{H}^{\rm int}_{RA}\mathcal{N}(\psi_{RA^{\prime}})].$ (62)
The following corollary directly follows from the above proposition.
Corollary 2.
For an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$, when the reference-output have noninteracting Hamiltonian $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes\widehat{H}_{A}+\widehat{H}_{R}\otimes\mathbbm{1}_{A}$ ($\widehat{H}^{\rm int}_{RA}=0$), then
$E[\mathcal{N}]=\sup_{\psi\in\operatorname{St}(A^{\prime})}\langle\widehat{H}_{A}\rangle_{\mathcal{N}(\psi)},$ (63)
where it suffices to consider optimization over all pure states $\psi_{A^{\prime}}$.
Theorem 5.
The thermodynamic free energy $F^{\beta}_{c}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is upper bounded as, when the reference $R$ is noninteracting with AA ($\widehat{H}_{RA}=\widehat{H}_{R}\otimes\mathbbm{1}_{A}+\mathbbm{1}_{R}\otimes\widehat{H}_{A}$),
$F^{\beta}_{c}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}].$ (64)
The bound is saturated for the replacer channels $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, for $\omega\in\operatorname{St}(A)$ and $\widehat{H}^{\rm int}_{RA}=0$, $F^{\beta}_{c}[\mathcal{R}^{\omega}]=E[\mathcal{R}^{\omega}]-\beta^{-1}S[\mathcal{R}^{\omega}]$.
Proof.
It follows from the definition of the free energy $F^{\beta}[\mathcal{N}]$ that
$F^{\beta}[\mathcal{N}]$
$=\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[E(\mathcal{N}(\psi_{RA^{\prime}}))-E(\psi_{R})-\beta^{-1}(S(\mathcal{N}(\psi_{RA^{\prime}}))\right$.
$\qquad\qquad\qquad\left.-S(\psi_{R}))+\beta^{-1}(\ln Z^{\beta}_{RA}-\ln Z^{\beta}_{R})\right],$
and $F^{\beta}[\mathcal{N}]=F^{\beta}_{c}[\mathcal{N}]+\beta^{-1}\ln Z^{\beta}_{A}$.
If $\widehat{H}_{RA}=\widehat{H}_{R}\otimes\mathbbm{1}_{A}+\mathbbm{1}_{R}\otimes\widehat{H}_{A}$, then $\ln Z^{\beta}_{RA}=\ln Z^{\beta}_{R}+\ln Z^{\beta}_{A}$, and
$F^{\beta}_{c}[\mathcal{N}]$ $=\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[E(\mathcal{N}(\psi_{RA^{\prime}}))-E(\psi_{R})\right$.
$\quad\qquad\qquad\left.-\beta^{-1}(S(\mathcal{N}(\psi_{RA^{\prime}}))-S(\psi_{R}))\right]$
$\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}].$ (65)
We conclude the proof by observing that for a replacer channel $\mathcal{R}^{\omega}$, $F^{\beta}_{c}[\mathcal{R}^{\omega}]=F^{\beta}_{c}(\omega)=E(\omega)-\beta^{-1}S(\omega)=E[\mathcal{R}^{\omega}]-\beta^{-1}S[\mathcal{R}^{\omega}]$, using the fact that $R$ and $A$ are noninteracting. ∎
Lemma 5.
The generalized resource-theoretic free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ and the generalized thermodynamic free energy $\mathbf{F}^{\beta}_{c}[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ upper bounds the generalized resource-theoretic free energy $\mathbf{F}^{\beta}(\Phi^{\mathcal{N}}_{RA})$ and the generalized thermodynamic free energy $\mathbf{F}^{\beta}_{c}(\Phi^{\mathcal{N}}_{RA})$ of its Choi state $\Phi^{\mathcal{N}}_{RA}$, respectively,
$\mathbf{F}^{\beta}[\mathcal{N}]$ $\geq\mathbf{F}^{\beta}(\Phi^{\mathcal{N}}_{RA}),$ (66)
$\mathbf{F}^{\beta}_{c}[\mathcal{N}]$ $\geq\mathbf{F}^{\beta}_{c}(\Phi^{\mathcal{N}}_{RA})+\beta^{-1}\ln|A^{\prime}|,$ (67)
when $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$ for $c\in\mathbbm{R}$.
Proof.
The proof follows from the observation that for $\widehat{H}_{R}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$, we have $\pi_{R}\otimes\gamma^{\beta}_{A}=\gamma^{\beta}_{R}\otimes\gamma^{\beta}_{A}=\gamma^{\beta}_{RA}=\frac{1}{|A^{\prime}|}\mathbbm{1}_{R}\otimes\mathbbm{1}_{A}=\frac{1}{|A^{\prime}|}\widehat{\gamma}_{RA}$ for $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$. ∎
VI.1 Extractable work from a quantum channel
[…]
$\delta E=\delta W+\delta Q.$ (68)
The entropy $S(\rho)$ of the system in thermodynamics is defined through its change due to the heat flow into the system $\delta Q_{\rm rev}$ during a reversible process at inverse temperature $\beta$,
$\delta S:=\beta\delta Q_{\rm rev}.$ (69)
A reversible, isothermal process is necessarily quasistatic and the system is always in thermodynamic equilibrium with the bath. Therefore, during a reversible process, the work done on the system is given by
$\delta W=\delta E-\beta^{-1}\delta S=\delta F_{c}^{\rm eq}.$ (70)
where $F_{c}^{\rm eq}(\rho_{\beta}):=E(\rho_{\beta})-\beta^{-1}S(\rho_{\beta})$ is the equilibrium free energy of the state $\rho_{\beta}$ that is in thermodynamical equilibrium with the bath. An example of such a state is the Gibbs state $\gamma_{\beta}=\frac{e^{-\beta\widehat{H}}}{Z^{\beta}}$. The equilibrium free energy is $F^{\rm eq}_{c}(\gamma_{\beta})=-\beta^{-1}\ln Z^{\beta}$.
Non-equilibrium thermodynamics.— If the system is not in equilibrium with the bath, the change $\delta S$ in the entropy of the system has contributions from the heat flow $\delta Q$ to the environment, which is called the reversible entropy production $\delta_{e}S=\beta\delta Q$, and also from an irreversible entropy production $\delta_{i}S$ due to change in correlations with the environment [77],
$\delta S=\beta\delta Q+\delta_{i}S.$ (71)
The work done during an irreversible process is given by
$\delta W$ $=\delta E-\delta Q$ (72)
$=\delta E-\beta^{-1}\delta S+\beta^{-1}\delta_{i}S$ (73)
$=\delta F^{\beta}_{c}+\beta^{-1}\delta_{i}S.$ (74)
where $F^{\beta}_{c}(\rho)=E(\rho)-\beta^{-1}S(\rho)$ is the non-equilibrium free energy, also called the thermodynamic free energy. Let us now consider the thermalization of a system with Hamiltonian $\widehat{H}$ in the non-equilibrium state $\rho$ to the Gibbs state $\gamma_{\beta}$. This transformation can be carried in two steps:
  1. 1.
    Quenching: We instantaneously change the Hamiltonian from $\widehat{H}$ to $\beta^{-1}\ln\rho$. With respect to this new Hamiltonian, the state $\rho$ is equilibrium state. The work done on the system during the quench is the change in the energy of the system, given by
[…]
    $)\displaystyle=\operatorname{tr}(\rho(-\beta^{-1}\ln\rho-\widehat{H}$)) (76)
    $=-\beta^{-1}\operatorname{tr}(\rho\ln\rho-\rho\ln\gamma_{\beta})+\beta^{-1}\ln Z^{\beta}$ (77)
    $=-F^{\beta}(\rho)-F_{c}^{\rm eq}(\gamma_{\beta})=-F_{c}^{\beta}(\rho).$ (78)
    where $F^{\beta}(\rho)=\beta^{-1}\operatorname{tr}(\rho\ln\rho-\rho\ln\gamma_{\beta})=\beta^{-1}D(\rho\|\gamma_{\beta})$ is the resource-theoretic free energy of the state $\rho$.
  2. 2.
    Reversible, isothermal process: We quasistatically change the Hamiltonian from $\beta^{-1}\ln Z^{\beta}$ back to $\widehat{H}$. Since the state at the initial point of this process is at equilibrium, the state at the final point of the reversible process will be $\gamma_{\beta}$. The work done during this reversible process is given by
    $\delta W_{\rm rev}$ $=\delta F^{\rm eq}_{c}=F^{\rm eq}_{c}(\gamma_{\beta})-F^{\rm eq}_{c}(\rho)=F^{\rm eq}_{c}(\gamma_{\beta}).$ (79)
The total work done on the system during this process is
$W_{\rho\to\gamma}$ $=\delta W_{\rm quench}+\delta W_{\rm rev}$ (80)
$=-F^{\beta}_{c}(\rho)+F_{c}^{\rm eq}(\gamma_{\beta})$ (81)
$=-F^{\beta}(\rho).$ (82)
This is also the minimum work that must be done on the system during the transformation $\rho\to\gamma^{\beta}$ [78]. Equivalently, the maximum work that can be extracted from an out-of-equilibrium state $\rho$ as it thermalizes to the Gibbs state $\gamma_{\beta}$ is given by
$W^{\rm ext}_{\rho\to\gamma_{\beta}}=F^{\beta}(\rho).$ (83)
Partial thermalization: We now consider a bipartite state $\psi_{RA^{\prime}}$ and a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ that acts locally on $A^{\prime}$, say $\rho_{RA^{\prime}}:=\mathcal{N}(\psi_{RA^{\prime}})$. We assume that the channel output and its reference are noninteracting, $\widehat{H}_{RA}=\widehat{H}_{R}\otimes\mathbbm{1}_{A}+\mathbbm{1}_{R}\otimes\widehat{H}_{A}=\widehat{H}_{R}+\widehat{H}_{A}$. We can extract work from this system $\rho_{RA}$ via partial thermalization to the state $\rho_{R}\otimes\gamma^{\beta}_{A}$ (equal to $\psi_{R}\otimes\rho_{A}$), with the following steps (see Fig. 1):
  1. 1.
    Decoupling: The state $\rho_{RA}$ is first decoupled to $\rho_{R}\otimes\rho_{A}$. In the task of decoupling, the information stored as correlation gets erased and only local information content on reduced states remain. This task of erasure releases heat [79], which can be used to extract useful work. The maximum extractable work from the decoupling of a bipartite state $\rho_{AB}$ in contact with a bath at inverse temperature $\beta$ is [80, 81],
    $W^{\rm ext}_{\rho_{RA}\to\rho_{R}\otimes\rho_{A}}=\beta^{-1}I(A;R)_{\rho_{RA}}.$ (84)
  2. 2.
    Partial quenching: We instantaneously change the Hamiltonian $\widehat{H}_{A}$ to $\beta^{-1}\ln\rho_{A}$ from which we can extract the work,
    $W^{\rm ext}_{\rm quench}=F_{c}^{\beta}(\rho_{A}).$ (85)
  3. 3.
    We perform reversible, isothermal driving of system A from the Hamiltonian $\beta^{-1}\ln\rho_{A}$ back to $\widehat{H}_{A}$. The extractable work during this process is given by
    $W^{\rm ext}_{\rm rev}=-F_{c}^{\rm eq}(\gamma^{\beta}_{A}).$ (86)
The total extractable work during the partial thermalization process, $\operatorname{id}_{R}\otimes\mathcal{N}(\psi_{RA^{\prime}})\to\operatorname{id}_{R}\otimes\mathcal{T}^{\beta}(\psi_{RA^{\prime}})$, is
$W^{\rm ext}_{\mathcal{N}(\psi_{RA^{\prime}})\to\mathcal{T}^{\beta}(\psi_{RA^{\prime}})}=\beta^{-1}I(A;R)_{\mathcal{N}(\psi_{RA^{\prime}})}+F^{\beta}(\mathcal{N}(\psi_{A^{\prime}})).$ (87)
We define the maximal extractable work $W^{\rm ext}_{\mathcal{N}\to\mathcal{T}^{\beta}}$ from a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ via partial thermalization with respect to the bath $\beta$ as
$W^{\rm ext}_{\mathcal{N}\to\mathcal{T}^{\beta}}:=\sup_{\psi\in\operatorname{St}(RA^{\prime})}W^{\rm ext}_{\mathcal{N}(\psi_{RA^{\prime}})\to\mathcal{T}^{\beta}(\psi_{RA^{\prime}})}.$ (88)
In the following theorem, we establish a direct relation between the resource-theoretic free energy $F^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}{A^{\prime}\to A}$ and its maximal extractable work via partial thermalization.
Theorem 6.
The resource-theoretic free energy $F^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}$ measures the maximal extractable work of the channel via partial thermalization,
$F^{\beta}[\mathcal{N}]=W^{\rm ext}_{\mathcal{N}\to\mathcal{T}^{\beta}}.$ (89)
The proof follows directly from Proposition 5. The theorem quantitatively establishes the direct relation of the resource-theoretic free energy of a channel with its thermodynamic utility in terms of extractable work.
VII Discussion
We develop an asymptotically reversible resource theory of athermality for quantum channels, using Gibbs-preserving superchannels as free operations. Table 1 summarizes a comparison of basic thermodynamic concepts for states (as known in the literature) and channels (all except entropy are introduced in this work). The (resource-theoretic and thermodynamic) free energy of a quantum channel quantitatively and operationally capture its distinguishability from the absolutely thermal channel with respect to a bath, thereby providing general and concrete form to the suggestions put forward in [19, 25, 31]. The discussions in [30, Section IV. A.] and [25, Sections V & VI] along with the findings in this work, provide a generic framework to extend the notion of the free energy of a quantum channel to the notion of the free energy of higher order quantum processes and the conditional free energy of a bipartite quantum processes. We will be exploring these directions rigorously and formally in future works.
Another future direction is to study a resource theory of athermality of quantum channels under other classes of physically and operationally motivated free operations, for example, superchannel generalization of thermal operations [12, 23]. The set of all superchannels that are superchannel generalization of thermal operations will be a subset of Gibbs preserving superchannels. We want to see if the associated resource theory will be reversible or irreversible in the light of [16]. Another future direction would be to inspect the behavior of free energy functions for quantum Markovian and non-Markovian dynamics and their implications on the capabilities of the dynamics to perform thermodynamic tasks [82, 83, 84, 85, 86].
Acknowledgements.
The authors thank Manabendra Nath Bera, Swati Choudhary, Karol Horodecki, and Uttam Singh for discussions. D.G.S. thanks the Department of Science and Technology, Government of India, for the INSPIRE fellowship. S.D. acknowledges support from the Science and Engineering Research Board (now ANRF), Department of Science and Technology, Government of India, under Grant No. SRG/2023/000217 and the Ministry of Electronics and Information Technology (MeitY), Government of India, under Grant No. 4(3)/2024-ITEA. S.D. also acknowledges support from the National Science Centre, Poland, grant Opus 25, 2023/49/B/ST2/02468. S.D. thanks the University of Gdańsk, Gdańsk, Poland for the hospitality during his visit.
Appendix A Detailed Proofs
For sanwiched Rényi relative entropy $D_{\alpha}$, $\alpha=1$ is used to denote (Umegaki) relative entropy as $\lim_{\alpha\to 1}D_{\alpha}(\cdot\|\cdot)=D(\cdot\|\cdot)$. We now prove results claimed in the main content.
A.1 Proof of Theorem 1
We begin the proof by proving two useful lemmas employed to prove the theorem.
Lemma 6.
Given two quantum channels $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{A^{\prime}\to A}$ such that $\frac{1}{2}\|\mathcal{N}-\mathcal{M}\|_{\diamond}\leq\varepsilon$, we have
$|F[\mathcal{N}]-F[\mathcal{M}]|\leq\frac{1}{\beta}(\varepsilon K+h_{2}(\varepsilon)),$ (90)
where $\varepsilon\in[0,1]$ and $h_{2}(\varepsilon)=-\varepsilon\ln(\varepsilon)-(1-\varepsilon)\ln(1-\varepsilon)$ is the binary Shannon entropy. $K$ is a positive real number such that
$K\geq\max\{D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}],D_{\infty}[\mathcal{K}\|\mathcal{T}^{\beta}]\}.$
Proof.
The condition $\frac{1}{2}\|\mathcal{N}-\mathcal{M}\|_{\diamond}\leq\varepsilon$ implies that $\frac{1}{2}\|\mathcal{N}(\psi_{RA^{\prime}})-\mathcal{M}(\psi_{RA^{\prime}})\|_{1}\leq\varepsilon$ for all $\psi_{RA^{\prime}}$. The continuity of relative entropy in the first argument has been tightened in [87, 88]. Here we use a slightly simpler bound given in [87, Theorem 1], from which it follows that
$D(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))-D(\mathcal{M}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$
$\leq\varepsilon K+h_{2}(\varepsilon),$ (91)
where
$K\geq D_{\infty}(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}})).$ (92)
is some positive real number. Therefore, we have
$D[\mathcal{N}\|\mathcal{T}^{\beta}]-D[\mathcal{M}\|\mathcal{T}^{\beta}]$
$\leq D(\mathcal{N}(\psi^{0}_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi^{0}_{RA^{\prime}}))-D(\mathcal{M}(\psi^{0}_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi^{0}_{RA^{\prime}}))$ (93)
$\leq\varepsilon K_{\mathcal{N}}+h_{2}(\varepsilon),$ (94)
where the state $\psi^{0}_{RA^{\prime}}$ maximizes $D(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$ and $K_{\mathcal{N}}\geq D_{\infty}(\mathcal{N}(\psi^{0}_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi^{0}_{RA^{\prime}}))=D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]$. Following the similar steps for the channel $\mathcal{M}$,
$D[\mathcal{M}\|\mathcal{T}^{\beta}]-D[\mathcal{N}\|\mathcal{T}^{\beta}]\leq\varepsilon K_{\mathcal{M}}+h_{2}(\varepsilon),$ (95)
where $K_{\mathcal{M}}\geq D_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]$. From Eqs. (A.1) and (95), we have
$|D[\mathcal{N}\|\mathcal{T}^{\beta}]-D[\mathcal{M}\|\mathcal{T}^{\beta}]|\leq\varepsilon K+h_{2}(\varepsilon)$ (96)
$\text{where,}\penalty 10000\ K=\max\{K_{\mathcal{N}},K_{\mathcal{M}}\}.$
The proof of the lemma follows by using the definition of free energy $F[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ in the above inequality. ∎
Lemma 7.
Given two quantum channels $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{A^{\prime}\to A}$ such that $\frac{1}{2}\|\mathcal{N}-\mathcal{M}\|_{\diamond}\leq\varepsilon$, we have
[…]
where $E_{\max}$ is the maximum eigenvalue of the Hamiltonian $\widehat{H}_{A}$.
Proof.
Using the properties of max-relative channel entropy, we have,
$D_{\infty}$ $[\mathcal{N}\|\mathcal{T}^{\beta}]-D_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]$ (98)
$=D_{\infty}(\Phi^{\mathcal{N}}\|\Phi^{\mathcal{T}^{\beta}})-D_{\infty}(\Phi^{\mathcal{M}}\|\Phi^{\mathcal{T}^{\beta}})$ (99)
$=D_{\infty}(\Phi^{\mathcal{N}}\|\pi_{R}\otimes\gamma^{\beta}_{A})-D_{\infty}(\Phi^{\mathcal{M}}\|\pi_{R}\otimes\gamma^{\beta}_{A}).$ (100)
The minimum eigenvalue of the second argument is given by
$\lambda_{0}:=\lambda_{\min}(\pi_{R}\otimes\gamma^{\beta}_{A})=\frac{\mathrm{e}^{-\beta E_{\max}}}{|A^{\prime}|Z^{\beta}}.$ (101)
where $E_{\max}$ is the largest eigenvalue of the Hamiltonian $\widehat{H}_{A}$. Using the bounds for uniform continuity of max-relative entropy given in [89], we have
$|D_{\infty}$ $[\mathcal{N}\|\mathcal{T}^{\beta}]-D_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]|$ (102)
$\leq\min\begin{cases}\ln(1+\varepsilon)+\ln\lambda_{0}^{-1},\\ \ln(1+\varepsilon\lambda_{0}^{-1}),\\ \ln((1+\varepsilon)(1+\varepsilon\lambda_{0}^{-1})-\varepsilon^{2}).\end{cases}$ (103)
Note that $Z^{\beta}\mathrm{e}^{\beta E_{\max}}>1$ for any Hamiltonian $\widehat{H}_{A}$. Therefore, $\lambda_{0}^{-1}$ is greater than $1$, which makes the second term in the above cases the minimum value for any Hamiltonian. Therefore, we have
$|D_{\infty}$ $[\mathcal{N}\|\mathcal{T}^{\beta}]-D_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]|\leq\ln(1+\varepsilon|A^{\prime}|Z^{\beta}\mathrm{e}^{\beta E_{\max}}).$ (104)
The proof follows from the definition of max-free energy. ∎
We now prove the theorem.
Theorem.
Let $\mathcal{N}_{A^{\prime}\to A}$ be an arbitrary quantum channel. The free energy $F^{\beta}[\mathcal{N}]$ satisfies all the thermodynamical (resource-theoretic) axiomatic properties (A1-A6).
For $\alpha\in(1,\infty)$, the sandwiched Rényi free energy $F_{\alpha}^{\beta}[\mathcal{N}]$ satisfies (A1-A5) and quasi-convexity, i.e., for quantum channel $\mathcal{N}=\sum_{x}p_{x}\mathcal{N}^{x}$ where $\{p_{x}\}_{x}$ is some probability distribution, $F_{\alpha}^{\beta}[\mathcal{N}]\leq\max_{x}F_{\alpha}^{\beta}[\mathcal{N}^{x}]$.
Proof.
The free energy of a channel satisfies all the axioms as shown below.
(A1) Monotonicity under Gibbs preserving superchannels: This follows from the monotonicity of sandwiched Rényi relative entropy between channels for $\alpha\in[\frac{1}{2},\infty)$ [90],
$D_{\alpha}[\mathcal{N}_{A^{\prime}\to A}\|\mathcal{T}^{\beta}_{A^{\prime}\to A}]$ $\geq D_{\alpha}[\Theta(\mathcal{N}_{A^{\prime}\to A})\|\Theta(\mathcal{T}^{\beta}_{A^{\prime}\to A})]$ (105)
$=D_{\alpha}[\Theta(\mathcal{N}_{A^{\prime}\to A})\|\mathcal{T}^{\beta}_{A^{\prime}\to A}],$ (106)
which imply $F_{\alpha}^{\beta}[\Theta(\mathcal{N}_{A^{\prime}\to A})]\leq F_{\alpha}^{\beta}[\mathcal{N}_{A^{\prime}\to A}]$. (A2) Reduction to states: For a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, for all $\alpha\in[\frac{1}{2},\infty)$, we have
$\beta F_{\alpha}[\mathcal{R}^{\omega}]$ $=D_{\alpha}[\mathcal{R}^{\omega}\|\mathcal{T}^{\beta}]$
$=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\alpha}(\mathcal{R}^{\omega}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$
$=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\alpha}(\psi_{R}\otimes\omega_{A}\|\psi_{R}\otimes\gamma^{\beta}_{A})$
$=D_{\alpha}(\omega_{A}\|\gamma^{\beta}_{A})=\beta F_{\alpha}(\omega).$ (107)
(A3) Minimum value at absolutely thermal channel: This follows from the fact that for $\alpha\in[\frac{1}{2},\infty)$, $D_{\alpha}(\rho\|\sigma)\geq 0$ if both $\rho$ and $\sigma$ are states, and $D_{\alpha}(\rho\|\sigma)=0$ if and only if $\rho=\sigma$. Therefore, given quantum channels $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{A^{\prime}\to A}$, $D_{\alpha}[\mathcal{N}\|\mathcal{M}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\alpha}(\mathcal{N}(\psi)\|\mathcal{M}(\psi))=0$ if and only if $\mathcal{N}(\varphi)=\mathcal{M}(\varphi)$ for all $\varphi\in\operatorname{St}(RA^{\prime})$ or $\mathcal{N}=\mathcal{M}$. Therefore, $F_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]\geq 0$ with equality if and only if $\mathcal{N}=\mathcal{T}^{\beta}$.
(A4) Continuity: We show the continuity of channel free energy $F^{\beta}[\mathcal{N}]$ in Lemma 6 and channel max-free energy $F^{\beta}_{\infty}[\mathcal{N}]$ in Lemma 7. The continuity of $F_{\alpha}^{\beta}[\mathcal{N}]$ for $\alpha\in[\frac{1}{2},\infty)$ can be shown similarly using the continuity of $D_{\alpha}(\rho\|\sigma)$ in the first argument, as given in [89].
(A5) Additivity: For two channels $\mathcal{N}_{A_{1}^{\prime}\to A_{1}}$ and $\mathcal{M}_{A_{2}^{\prime}\to A_{2}}$ such that the output interaction Hamiltonian $\widehat{H}_{A_{1}A_{2}}$ is zero, the absolutely thermal channel can be written as $\mathcal{T}^{\beta}_{A_{1}^{\prime}A_{2}^{\prime}\to A_{1}A_{2}}=\mathcal{T}^{\beta}_{A_{1}^{\prime}\to A_{1}}\otimes\mathcal{T}^{\beta}_{A_{2}^{\prime}\to A_{2}}=:\mathcal{T}^{\beta}_{1}\otimes\mathcal{T}^{\beta}_{2}$. In this case, the additivity of the generalized free energy of channel follows from the additivity of the generalized channel divergence since
$\beta F_{\alpha}[\mathcal{N}\otimes\mathcal{M}]=D_{\alpha}[\mathcal{N}\otimes\mathcal{M}\|\mathcal{T}^{\beta}_{1}\otimes\mathcal{T}^{\beta}_{2}].$ (108)
We will show that the sandwiched Rényi channel divergence is additive under tensor product with respect to the absolutely thermal channel, i.e.
$D_{\alpha}[\mathcal{N}\otimes\mathcal{M}\|\mathcal{T}^{\beta}_{1}\otimes\mathcal{T}^{\beta}_{2}]$ (109)
$=D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}_{1}]+D_{\alpha}[\mathcal{M}\|\mathcal{T}^{\beta}_{2}].$ (110)
The proof of the inequality $\geq$” follows from the definition of the sandwiched Rényi channel divergence,
$D_{\alpha}[\mathcal{N}\otimes\mathcal{M}\|\mathcal{T}^{\beta}_{1}\otimes\mathcal{T}^{\beta}_{2}]$
$=\sup_{\psi\in\operatorname{St}(RA_{1}^{\prime}A_{2}^{\prime})}D_{\alpha}(\mathcal{N}\otimes\mathcal{M}(\psi)\|\mathcal{T}^{\beta}_{1}\otimes\mathcal{T}^{\beta}_{2}(\psi))$ (111)
$\geq\sup_{\varphi\in\operatorname{St}(R_{1}A_{1}^{\prime})\otimes\operatorname{St}(R_{2}A_{2}^{\prime})}D_{\alpha}(\mathcal{N}\otimes\mathcal{M}(\varphi)\|\mathcal{T}^{\beta}_{1}\otimes\mathcal{T}^{\beta}_{2}(\varphi))$ (112)
$=D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}_{1}]+D_{\alpha}[\mathcal{M}\|\mathcal{T}^{\beta}_{2}],$ (113)
where we have used the additivity of the sandwiched Rényi divergence of states under the tensor product for $\alpha\in[\frac{1}{2},\infty)$.
To prove the inequality $\leq$”, we use the following result from [64, Lemma 38 and Proposition 41], for all $\alpha\in[1,\infty)$, for an arbitrary state $\rho_{RA^{\prime}}$, a channel $\mathcal{N}_{A^{\prime}\to A}$ , and a replacer channel $\mathcal{R}_{A^{\prime}\to A}$ which replaces the state with a state $\tau$,
$D_{\alpha}(\mathcal{N}(\rho_{RA^{\prime}})\|\sigma_{R}\otimes\tau_{A})$ (114)
$\leq D_{\alpha}(\mathcal{N}\|\mathcal{R}^{\tau})+D_{\alpha}(\rho_{RA^{\prime}}\|\sigma_{RA^{\prime}}).$ (115)
Let $\psi_{RA_{1}^{\prime}A_{2}^{\prime}}$ be an arbitrary pure state and states $\rho_{R^{\prime}A_{1}^{\prime}}$ and $\sigma_{R^{\prime}A_{1}^{\prime}}$ be such that $R^{\prime}=RA_{2}$, i.e., $\mathcal{M}(\psi_{RA_{1}^{\prime}A_{2}^{\prime}})=\rho_{R^{\prime}A_{1}^{\prime}}$ and $\mathcal{T}^{\beta}_{A_{2}^{\prime}\to A_{2}}(\psi_{RA_{1}^{\prime}A_{2}^{\prime}})=\sigma_{R^{\prime}A_{1}^{\prime}}$,
$D_{\alpha}(\mathcal{N}\otimes\mathcal{M}(\psi_{RA_{1}^{\prime}A_{2}^{\prime}})\|\mathcal{T}^{\beta}_{A_{1}^{\prime}\to A_{1}}\otimes\mathcal{T}^{\beta}_{A_{2}^{\prime}\to A_{2}}(\psi_{RA_{1}^{\prime}A_{2}^{\prime}}))$ (116)
$\leq D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]+D_{\alpha}(\mathcal{M}(\psi_{RA_{1}^{\prime}A_{2}^{\prime}})\|\mathcal{T}^{\beta}_{A_{2}^{\prime}\to A_{2}}(\psi_{RA_{1}^{\prime}A_{2}^{\prime}}))$ (117)
$\leq D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]+D_{\alpha}[\mathcal{M}\|\mathcal{T}^{\beta}].$ (118)
This implies the additivity of $F_{\alpha}^{\beta}[\mathcal{N}]$ for $\alpha\in[1,\infty)$.
(A6) Joint quasi-convexity: It is known that the sandwiched Rényi divergence is jointly convex for all $\alpha\in[\frac{1}{2},1]$ and jointly quasi-convex for all $\alpha\in(1,\infty)$. For $\alpha\in[\frac{1}{2},1]$, we have
$D_{\alpha}$ $\left[\sum_{x}p_{x}\mathcal{N}^{x}\|\mathcal{T}^{\beta}\right]$
$=\sup_{\psi_{RA^{\prime}}}D_{\alpha}\left(\sum_{x}p_{x}\mathcal{N}^{x}(\psi)\|\mathcal{T}^{\beta}(\psi)\right)$ (119)
$\leq\sup_{\psi_{RA^{\prime}}}\sum_{x}p_{x}D_{\alpha}(\mathcal{N}^{x}(\psi)\|\mathcal{T}^{\beta}(\psi)).$ (120)
Using the definition of free energy of channels, we have
$F_{\alpha}^{\beta}\left[\sum_{x}p_{x}\mathcal{N}^{x}\right]\leq\sum_{x}p_{x}F_{\alpha}^{\beta}[\mathcal{N}^{x}]\penalty 10000\ \text{for}\penalty 10000\ \alpha\in\left[\frac{1}{2},1\right].$ (121)
Similarly, using the quasi-convexity of sandwiched Rényi divergence for $\alpha\in(1,\infty)$, we have the quasi-convexity for free energy,
$F_{\alpha}\left[\sum_{x}p_{x}\mathcal{N}^{x}\right]\leq\max_{x}F_{\alpha}[\mathcal{N}^{x}].$ (122)
A.2 Proof of Lemma 1
Lemma.
The sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]$, $\alpha\in[\frac{1}{2},\infty)$, is nonincreasing under the action of Gibbs-subpreserving superchannel. It remains invariant under the action of Gibbs preserving unitary superchannels.
Proof.
Given the channel $\mathcal{N}_{A^{\prime}\to A}$ and absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$, the sandwiched Rényi relative entropy of a channel is monotonically non-increasing under the action of a superchannel $\Theta\in\operatorname{SCh}((A^{\prime},A),(B^{\prime},B))$ [90]. i.e. for all $\alpha\in[\frac{1}{2},\infty)$,
$D_{\alpha}[\Theta(\mathcal{N})\|\Theta(\mathcal{T}^{\beta})]\leq D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}].$ (123)
For positive semidefinite operators $\rho\neq 0$, $\sigma$ and $\sigma^{\prime}$, if $\sigma^{\prime}\geq\sigma$, we have $D_{\alpha}(\rho\|\sigma^{\prime})\leq D_{\alpha}(\rho\|\sigma)$ for $\alpha\in[\frac{1}{2},\infty)$ ([91, Proposition 4]). Therefore, under the action of a $\mathcal{T}^{\beta}$-subpreserving superchannel, we have
$D_{\alpha}(\Theta(\mathcal{N})(\rho_{RB^{\prime}})\|\mathcal{T}^{\beta}_{B^{\prime}\to B}(\rho_{RB^{\prime}}))$
$\leq D_{\alpha}(\Theta(\mathcal{N})(\rho_{RB^{\prime}})\|\Theta(\mathcal{T}^{\beta}_{A^{\prime}\to A})(\rho_{RB^{\prime}})),$
for all states $\rho_{RB^{\prime}}$. Taking the supremum over all states on both sides,
$D_{\alpha}[\Theta(\mathcal{N})\|\mathcal{T}^{\beta}_{B^{\prime}\to B}]\leq D_{\alpha}[\Theta(\mathcal{N})\|\Theta(\mathcal{T}^{\beta}_{A^{\prime}\to A})].$ (124)
Using Eqs. (123) and (124), we have
$D_{\alpha}[\Theta(\mathcal{N})\|\mathcal{T}^{\beta}_{B^{\prime}\to B}]\leq D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}_{A^{\prime}\to A}],$ (125)
for all $\alpha\in[\frac{1}{2},\infty)$. This implies the first part of the lemma.
The invariance of $F_{\alpha}^{\beta}[\mathcal{N}]$ for $\alpha\in[\frac{1}{2},\infty)$ under Gibbs preserving unitary superchannel follows directly from [30, Proposition 1]. ∎
A.3 Proof of Theorem 2
Lemma 8.
Given a state $|\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$ where $\psi_{i}$ are pure states, the max-relative entropy is bounded as follows
[…]
Proof.
The max-relative entropy $D_{\infty}(\rho\|\sigma)$ can be written as
[…]
$=\ln\|AA^{\dagger}\|_{\infty}$ (129)
$=\ln\lambda_{\max}(AA^{\dagger}).$ (130)
[…]
$\max_{i}(G_{ii})\leq\lambda_{\max}(G)\leq\operatorname{tr}(G).$ (131)
[…]
Theorem.
For $\alpha\in\{1,\infty\}$ and $|A^{\prime}|=|A|$, the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is maximum if and only if $\mathcal{N}$ is a unitary channel. For all unitary channels $\mathcal{U}_{A^{\prime}\to A}$, $F^{\beta}_{\alpha}[\mathcal{U}]=F^{\beta}_{\alpha}[\operatorname{id}_{A^{\prime}\to A}]$.
Proof.
Given a state $\rho_{RA}$ and its decomposition into pure states $\rho_{RA}=\sum_{x}p_{x}\psi^{x}_{AB}$, we use the joint quasi-convexity of the sandwiched Rényi entropy to observe that for $\alpha\in[1,\infty)$
$D_{\alpha}(\sum_{x}p_{x}\psi^{x}_{AB}\|\sum_{x}p_{x}\rho^{x}_{R}\otimes\gamma^{\beta}_{A})\leq\max_{x}D_{\alpha}(\psi^{x}_{RA}\|\rho^{x}_{R}\otimes\gamma^{\beta}_{A}).$ (132)
where $\rho^{x}_{R}=\operatorname{tr}_{A}\psi_{RA}^{x}$. Therefore, $D_{\alpha}(\rho_{RA}\|\rho_{R}\otimes\gamma_{A})$ achieves the maximum value, for $\alpha\in[1,\infty)$, if $\rho_{RA}$ is a pure state. This implies that $D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]$ achieves the maximum value if $\mathcal{N}$ is a unitary channel.
We will show that the ”only if” condition is satisfied for $\alpha=1$ and $\infty$ by proving that the equality in Eq. (132) holds if and only if $\rho_{RA}$ is a pure state.
For $\alpha=1$, relative entropy is jointly convex, and therefore for $\rho=\sum_{i}p_{i}\psi_{i}$, where $\psi_{i}$ are pure, we have
$D(\rho\|\sigma)\leq\sum_{i}p_{i}D(\psi_{i}\|\sigma).$ (133)
The equality in the above holds if and only if $\rho$ is pure [92].
For the case of $\alpha\to\infty$ Let the equality for the quasi-convexity condition be satisfied, i.e. $D_{\infty}(\rho\|\sigma)=\max_{i}D_{\infty}(\psi_{i}\|\sigma)$, then using Lemma 8
[…]
Since the weighted average cannot be greater than the largest element of the set, the above condition is satisfied if and only if $\rho$ is a pure state.
A.4 Proof of Theorem 3
Theorem.
For any error $\varepsilon\in[0,1]$ and a given resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, the single-shot athermality distillation and formation are proportional to the $\varepsilon$-hypothesis-testing free energy and the $\varepsilon$-max-free energy of the channel $\mathcal{N}$, respectively,
$\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ $=\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (136)
$\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ $=\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (137)
$F^{\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
Proof.
A Gibbs preserving superchannel (GPSC) $\Theta_{\psi}^{\Lambda}$, defined through a pure state $\psi$ and an effect operator (element of a POVM) $0\leq\Lambda\leq\mathbbm{1}_{A}$, that acts on the resource $(\mathcal{N},\mathcal{T}^{\beta})$ to yield $(\operatorname{id}_{m},\mathcal{R}^{\pi})$ can be written as the following preparation,
$\Theta^{\Lambda}_{\psi}(\mathcal{N})=\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\operatorname{id}+\operatorname{tr}(\mathcal{N}(\psi)(\mathbbm{1}_{A}-\Lambda))\frac{1}{m^{2}-1}\operatorname{id}^{\perp}.$ (138)
The Gibbs preserving constraint on the superchannel, i.e. $\Theta_{\psi}^{\Lambda}(\mathcal{T}^{\beta})=\mathcal{R}^{\pi}$, can be written as $\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda)=\frac{1}{m^{2}}$, and the conversion distance $d_{\Theta_{\psi}^{\Lambda}}((\mathcal{N},\mathcal{T}^{\beta})\rightarrow(\operatorname{id},\mathcal{R}^{\pi}))$ can be given by $1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)$. Recalling the definition of the $\varepsilon$-single-shot distillation, we have the following:
$\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$
$=\sup_{m,\Theta}\{\ln m:\penalty 10000\ d_{\Theta_{\psi}^{\Lambda}}((\mathcal{N},\mathcal{T}^{\beta})\rightarrow(\operatorname{id},\mathcal{R}^{\pi}))\leq\varepsilon\}$ (139)
$=\sup_{m,\Lambda,\psi}\{\ln m:\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon,\penalty 10000\ \operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda)=\frac{1}{m^{2}}\}$ (140)
$=\sup_{\psi,\Lambda}\{-\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon\}$ (141)
$=\sup_{\psi}-\inf_{\Lambda}\{\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon\}$ (142)
$=\sup_{\psi}\frac{1}{2}D_{H}^{\varepsilon}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))$ (143)
$=\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}].$ (144)
Similarly, the $\varepsilon$-single-shot cost can be written as
$\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon;\Theta(\mathcal{R}^{\pi})=R^{\gamma}\}$ (145)
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon;$
$\hskip 56.9055ptR^{\gamma}=\frac{1}{m}\Theta(\operatorname{id})+\frac{1}{m}\sum_{i}\Theta(\mathcal{W}_{i})\}$ (146)
$=\inf_{\Theta}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon;m^{2}R^{\gamma}\geq\Theta(\operatorname{id})\}.$ (147)
Replacing $\mathcal{E}=\Theta(\operatorname{id})$ in the above relation, we have
$\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{)}$
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon;m^{2}R^{\gamma}\geq\mathcal{E}\}$ (148)
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon;2\ln m\geq D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]\}$ (149)
$=\inf_{m}\{\frac{1}{2}D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]:\penalty 10000\ \|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon\}$ (150)
$=\frac{1}{2}D^{\varepsilon}_{\max}[\mathcal{N}\|\mathcal{T}^{\beta}].$ (151)
A.5 SDP for the max-free energy
The max-free energy $F_{\infty}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is half of the logarithm of the optimal value of the following semidefinite program (SDP) (strong duality):
$\underline{\textit{Primal}}:$
$\begin{cases}\text{minimize}&\lambda\in\mathbbm{R}\\[6.0pt] \text{subject to}&\Phi^{\mathcal{N}}_{RA}\leq\lambda(\pi_{R}\otimes\gamma_{A}^{\beta}),\\[6.0pt] &\lambda\geq 0.\end{cases}$
$\underline{\textit{Dual}}:$
$\begin{cases}\text{maximize}&\operatorname{tr}(\Phi^{\mathcal{N}}_{RA}X_{RA})\\[6.0pt] \text{subject to}&\operatorname{tr}((\pi_{A}\otimes\gamma_{A}^{\beta})X_{RA})\leq 1,\\[6.0pt] &X_{RA}\geq 0.\end{cases}$
It follows from the fact that the max-relative entropy between channels is an SDP where strong duality holds [48], i.e., optimal solution of the primal and dual problems are the same.
A.6 Proof of Proposition 3
Proposition.
The max-free energy $F^{\beta}_{\infty}[\mathcal{U}]$ of a unitary quantum channel $\mathcal{U}_{A^{\prime}\to A}$ is
$F^{\beta}_{\infty}[\mathcal{U}]=\beta^{-1}\ln\operatorname{tr}\left[{(\gamma^{\beta}_{A})}^{-1}\right].$ (152)
If the Hamiltonian of $A$ is trivial or in general $\widehat{H}_{A}=c\mathbbm{1}_{A}$ for $c\in\mathbbm{R}$, then $F^{\beta}_{\infty}[\mathcal{U}]=2\beta^{-1}\ln|A|$.
Proof.
Using the property of the max-relative entropy of quantum channels
$\beta F_{\infty}^{\beta}[\mathcal{U}]=D_{\infty}[\mathcal{U}\|\mathcal{T}^{\beta}]=D_{\infty}(\Phi^{\mathcal{U}}\|\Phi^{\mathcal{T}^{\beta}}),$ (153)
[…]
$\beta F_{\infty}^{\beta}[\mathcal{U}]$ $=D_{\infty}(\Phi^{\mathcal{U}}\|\Phi^{\mathcal{T}})$ (154)
[…]
$=\ln\left(\operatorname{tr}\left[\gamma_{\beta}^{-1}\right]\right).$ (155)
If $\widehat{H}_{A}=c\mathbbm{1}_{A}$, $\operatorname{tr}\left[(\gamma^{\beta}_{A})^{-1}\right]=\operatorname{tr}(|A|\mathbbm{1}_{A})=|A|^{2}$ and $F^{\beta}_{\infty}[\mathcal{U}]=2\beta^{-1}\ln|A|$. ∎
A.7 Proof of Proposition 4
Proposition.
For a quantum channel $\mathcal{N}$ and $\varepsilon\in(0,1)$ the athermality distillation and formation of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ satisfy the following trade-off relation,
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{1-\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right).$ (156)
Proof.
For $\varepsilon\in(0,1)$ and $\rho\in\operatorname{St}(A)$, $\sigma\in\operatorname{Pos}(A)$, we have [93, 57]
$D^{\sqrt{\varepsilon}}_{\infty}(\rho\|\sigma)$ $\leq D^{1-\varepsilon}_{\mathrm{H}}(\rho\|\sigma)+\ln\left(\frac{1}{1-\varepsilon}\right).$ (157)
Note that given a channel $\mathcal{M}_{A^{\prime}\to A}$ s.t. $\mathcal{M}\in\mathcal{B}^{\epsilon}[\mathcal{N}]$ implies that $\mathcal{M}(\psi)\in\mathcal{B}^{\varepsilon}(\mathcal{N}(\psi))$ for all $\psi\in\operatorname{St}(RA^{\prime})$. Therefore, the smoothed max-relative channel entropy between any channel $\mathcal{N}$ and $\mathcal{P}$ satisfies the following inequality [cf. 38, Lemma 1],
$D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{P}]$ $=\sup_{\psi\in\operatorname{St}(RA^{\prime})}\sup_{\mathcal{M}\in\mathcal{B}^{\varepsilon}[\mathcal{N}]}D(\mathcal{M}(\psi)\|\mathcal{P}(\psi))$ (158)
$\leq\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\infty}^{\varepsilon}(\mathcal{N}(\psi)\|\mathcal{P}(\psi)).$ (159)
Using the above relation, we have
$F_{\infty}^{\beta,\sqrt{\varepsilon}}[\mathcal{N}]$
$=\frac{1}{2\beta}D_{\infty}^{\sqrt{\varepsilon}}[\mathcal{N}\|\mathcal{T}^{\beta}]$ (160)
$\leq\frac{1}{2\beta}\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\infty}^{\sqrt{\varepsilon}}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))$ (161)
$\leq\frac{1}{2\beta}\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\mathrm{H}}^{1-\varepsilon}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right)$ (162)
$=\frac{1}{2\beta}D_{\mathrm{H}}^{1-\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right)$ (163)
$=\frac{1}{2}F_{\mathrm{H}}^{\beta,1-\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right).$ (164)
A.8 Free energy of a golden unit
Lemma 9.
Given a unitary channel $\mathcal{U}_{A^{\prime}\to A}$, the minimum value of its max-free energy $F^{\beta}_{\infty}[\mathcal{U}]$ is achieved for the output with trivial Hamiltonian, $\widehat{H}_{A}=c\mathbbm{1}_{A}$. In particular,
$\min_{\widehat{H}_{A}}F^{\beta}_{\infty}[\mathcal{U}]=2\beta^{-1}\ln|A|.$ (165)
Proof.
From Proposition 3 we have
$\beta F_{\infty}^{\beta}[\mathcal{U}]$ $=\ln\operatorname{tr}(\gamma_{\beta}^{-1})$ (166)
$=\ln\operatorname{tr}{\operatorname{e}}^{\beta\widehat{H}_{A}}+\ln\operatorname{tr}{\operatorname{e}}^{-\beta\widehat{H}_{A}}.$ (167)
We analyze the vanishing conditions of the directinanl derivative with respect to $\widehat{H}_{A}$, that is $\nabla_{\widehat{H}_{A}}F_{\infty}^{\beta}[\mathcal{U}]=0$, which implies that
$\frac{{\operatorname{e}}^{\beta\widehat{H}_{A}}}{\operatorname{tr}({\operatorname{e}}^{\beta\widehat{H}_{A}})}-\frac{{\operatorname{e}}^{-\beta\widehat{H}_{A}}}{\operatorname{tr}({\operatorname{e}}^{-\beta\widehat{H}_{A}})}=0.$ (168)
For all eigenvalues $E_{i}$ of $\widehat{H}_{A}$, we have
$\frac{{\operatorname{e}}^{\beta E_{i}}}{\operatorname{tr}({\operatorname{e}}^{\beta\widehat{H}_{A}})}$ $=\frac{{\operatorname{e}}^{-\beta E_{i}}}{\operatorname{tr}({\operatorname{e}}^{-\beta\widehat{H}_{A}})}$ (169)
$\text{or},\penalty 10000\ {\operatorname{e}}^{2\beta E_{i}}$ $=\frac{\operatorname{tr}({\operatorname{e}}^{\beta\widehat{H}_{A}})}{\operatorname{tr}({\operatorname{e}}^{-\beta\widehat{H}_{A}})},\penalty 10000\ \forall i.$ (170)
Since the right-hand side is independent of $i$, we conclude that all the eigenvalues of $\widehat{H}_{A}$ must be the same, which corresponds to the Hamiltonian of the form $\widehat{H}_{A}=c\mathbbm{1}_{A}$. For such Hamiltonian, we have $F_{\infty}^{\beta}[\mathcal{U}]=2\beta^{-1}\ln|A|$.
To see that this value is the global minima for $F_{\infty}^{\beta}[\mathcal{U}]$ note that
$2^{\beta F_{\infty}^{\beta}[\mathcal{U}]}$ $=\operatorname{tr}{\operatorname{e}}^{\beta\widehat{H}_{A}}\operatorname{tr}^{-\beta\widehat{H}_{A}}$ (171)
$\geq(\operatorname{tr}{\operatorname{e}}^{\frac{1}{2}\beta\widehat{H}_{A}}{\operatorname{e}}^{-\frac{1}{2}\beta\widehat{H}_{A}})^{2}$ (172)
$=(\operatorname{tr}\mathbbm{1}_{A})^{2}$ (173)
$=|A|^{2}.$ (174)
Here we have used the Cauchy-Schwarz inequality for Hilbert-Schmidt inner product: $(\operatorname{tr}(A^{\dagger}B))^{2}\leq\operatorname{tr}(A^{\dagger}A)\operatorname{tr}(B^{\dagger}B)$. We conclude that $F_{\infty}^{\beta}[\mathcal{U}]\geq 2\beta^{-1}\ln|A|$ and the output Hamiltonian $\widehat{H}_{A}=c\mathbbm{1}_{A}$ gives the minimum free energy for unitary channels. ∎